### v1

# FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

Abstract
Researchers have proposed to use data of human preference feedback to fine-tune text-to-image generative models. However, the scalability of human feedback collection has been limited by its reliance on manual annotation. Therefore, we develop and test a method to automatically annotate user preferences from their spontaneous facial expression reaction to the generated images. We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images. Specifically, AU4 (brow lowerer) is most consistently reflective of negative evaluations of the generated image. This can be useful in two ways. Firstly, we can automatically annotate user preferences between image pairs with substantial difference in AU4 responses to them with an accuracy significantly outperforming state-of-the-art scoring models. Secondly, directly integrating the AU4 responses with the scoring models improves their consistency with human preferences. Additionally, the AU4 response best reflects the user’s evaluation of the image fidelity, making it complementary to the state-of-the-art scoring models, which are generally better at reflecting image-text alignment. Finally, this method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks. The code is available at https://github.com/ShuangquanFeng/FERGI, and the dataset is also available at the same link for research purposes.
1 Introduction
The rapid recent advancements in text-to-image generative models have enabled the generation of high-fidelity images aligned with text prompts [50, 56, 59, 57, 58, 54]. To better improve the fidelity of the generated images and their alignment with the text prompts, researchers collected large-scale datasets of human preference feedback on images generated by text-to-image models [73, 39, 32, 72, 71, 80] and proposed methods of fine-tuning the models with data of human preference feedback [73, 38, 21].
However, the scalability of human feedback collection has still been limited by its reliance on manual annotation. Therefore, we develop and test a system to automatically annotate user preferences for text-to-image generation from their spontaneous facial expression reaction to the generated images. This automatic annotation requires zero additional effort from the real users of text-to-image generators and is thus expected to increase the scalability of human preference feedback data collection.
We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion v1.4 [58] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input. We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [19]), in the facial expression reaction videos with a pre-trained model and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images. We further propose a method of automatically annotating user preferences by detecting the activation of AU4 in the facial expression reaction to the generated images and show that this evaluation metric best reflects the user’s evaluation of the image fidelity, making it complementary to the state-of-the-art scoring models, which are generally better at reflecting image-text alignment.
2 FERGI Dataset
2.1 Participants
39 participants were recruited from the SONA system of University of California San Diego (UCSD) and completed the study asynchronously on personal computers. 6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not permitting sharing of their video recordings for research. Therefore, we ended up with a dataset of 33 participants that is available upon request for research purposes at https://github.com/ShuangquanFeng/FERGI. The collection and usage of human participant data in the dataset were approved by the Institutional Review Board (IRB) of UCSD.
2.2 Data Collection Procedure
Each participant completes multiple sessions of data collection. In each session, the participant drafts one text prompt and views 5 images generated from the input text prompt. A flow chart of the procedure of each session is shown in Fig. 1. Details are explained below.
2.2.1 Prompt Creation
There are two types of sessions, structured input and free-form input. In both types, the participant can freely draft any text prompt as the input as long as it is not NSFW. The only difference between the two types lies in how the input text prompt is created. In a structured input session, the participant fills out a form to separately specify different elements desired in the images to generate: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords. Based on the contents of the form, a text prompt is then automatically generated by the Large Language Model (LLM) OpenAI gpt3.5-turbo. The participant can freely adjust the generated prompt as the final input text prompt. In a free-form input, the participant directly enters the entire final input text prompt.
2.2.2 Recording
After the input text prompt is finalized, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately. The webcam is then used to record throughout the rest of the session. The configuration of the raw videos vary based on the participant’s device and browser. Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos are standardized to 30 FPS afterwards.
2.2.3 Image Presentation, Annotation, and Ranking
In each session, a total of 5 images generated by Stable Diffusion v1.4 [58] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images). The image presentation and annotation follows the webcam preview. The detailed procedures for the presentation and annotation of each generated image are as follows:
  • Baseline (5 seconds): The website displays the “preparing image” status text at the center of the page along with the input text prompt. No specific facial-expression-eliciting event is expected to happen during this period.
  • Present (5 seconds): The generated image is presented at the center of the webpage along with the input text prompt. The participant’s facial expression reaction to the generated image is collected within this interval.
  • Annotate (no time restriction): The website displays the prompt, the generated image (moved to the right side), and the annotation survey. The participant annotates the image manually by filling out the survey during this interval. The survey was adapted from the survey designed for ImageReward [73] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (which was fixed afterwards) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis. (multi-choice question), and emotions felt when seeing the image (multi-choice question).
After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server.
2.3 Data Preprocessing
For each image, two clips are extracted from the recorded video: a 5-second “baseline clip” of the participant’s facial expression during the baseline period and a 5-second “reaction clip” of the participant’s facial expression during the image presentation period.
Stable Diffusion v1.4 [58] outputs an all-black image when it detects the original output image might be inappropriate. A total of 53 such images out of 2880 total generated were generated in the dataset. The data associated with these images were excluded from further analysis, resulting in less than 5 valid images for some input text prompts.
Figure 1: Procedure of one session in data collection.
3 AU Model Training
We analyze the facial expression reaction of the users by estimating the activation of their AUs. We trained our AU estimation model on the DISFA [49, 48] and DISFA+ [47] datasets. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). Video frames are first preprocessed with face detection [44], face alignment [6], and a combination of histogram equalization and linear mapping [35] before being fed into the model for training, testing, or inference. For model training, we use the neural network IR-50 [11] pre-trained on Glink360k [1] and fine-tune it on the DISFA and DISFA+ datasets [49, 47]. A single network learns the estimation of all AUs labeled in the datasets with both regression and ordinal classification [51].
4 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips.
4.1 Data Filtering
Robustness against face occlusion and faces with non-frontal poses is an ongoing challenge for facial expression recognition models [79, 27]. Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and has not been specifically adapted for non-frontal faces or faces with occlusion, so it is not suitable for estimating AUs of such faces. We address this issue by excluding such data from analysis. Specifically, for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [44]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [6]). Then, if a clip has more than $20\%$ of its total frames being excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
4.2 AU Activation Value
We compute an AU activation value $\alpha_{i}$ for each trained AU (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
  • Firstly, we estimate the intensity of each AU for each frame of the video clip with the AU model: $\hat{y}_{i,\mathrm{reg}}^{(1)},\hat{y}_{i,\mathrm{reg}}^{(2)},\dots$, where the superscript indicates the index of the frame.
  • Secondly, we compute the moving window mean of the estimated AU intensities for every 0.1 seconds (3 frames for 30 FPS):
    $\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)}=\frac{1}{3}\Sigma_{k^{\prime}=k}^{k+2}\hat{y}_{i,\mathrm{reg}}^{(k^{\prime})}$ (1)
    for all $k$ with defined $\hat{y}_{i,\mathrm{reg}}^{(k+1)}$ and $\hat{y}_{i,\mathrm{reg}}^{(k+2)}$ (as they may be undefined for frames at the end of the video clip or frames adjacent to excluded frames).
  • Finally, we compute the AU activation of the clip as the difference between the maximum AU intensity within the video clip and the AU intensity of the first 0.1 seconds:
    $\alpha_{i}=\max_{k}(\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)})-\bar{\hat{y}}_{i,\mathrm{reg}}^{(1)}.$ (2)
5 Experiments
5.1 Statistical Analysis
(a) AU4
(b) AU9
(c) AU1
(d) AU2
(e) AU5
(f) AU12
(g) AU20
[…]
(a) disappointed
(b) satisfied
(c) surprised
(d) disgusted
(e) amused
(f) scared
Figure 3: Reported emotions highly correlated with AU activation values. The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?” Each subfigure shows the results for a different emotion (indicated in the subcaptions) and contains multiple pairs of boxplots representing the results for different AUs. The x-axis represents the indices of the AUs while the y-axis represents the AU activation values in the reaction clip of each image. The parenthesized numbers in the legends represent the number of responses of that type. In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion. In each boxplot, the bottom of the box represents the first quartile (25th percentile) of the AU activation values, the top of the box represents the third quartile (75th percentile), and the line in the middle of the box represents the median.
We started with analyzing statistical relationships between the AU activation values and answers in the annotation survey for each image. The primary results are as follows:
  • Overall Ratings. The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs. 1(a) and 1(b) (Spearman correlation: $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9).
[…]
  • Emotions. For the question “Did you feel any of the following emotions when you saw the image?”, the participant can decide whether to include the six options (disappointed, satisfied, surprised, disgusted, amused, scared) in the answer independently (they can select 0 to 6 of the options as their answer). We found that whether they reportedly feel each of the emotions are significantly correlated with the activation values of multiple AUs.
    • Disappointed. ”Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig. 2(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. ”Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig. 2(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
    • Surprised. ”Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig. 2(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
    • Disgusted. ”Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig. 2(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
    • Amused. ”Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig. 2(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
    • Scared. ”Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig. 2(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$, with 12 representing the 12 investigated AUs, 2 representing the two tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 representing the 6 tests for 6 emotions.
5.2 Automatic Annotation of User Preferences between Image Pairs Using AU4
We have observed that AU4 activation value has the highest and most consistent (negative) correlation with the raw overall ratings in the previous subsection. Therefore, we use the participants’ AU4 activation values to automatically annotate their preferences between images generated based on the same input text prompt.
5.2.1 Participant Exclusion
Multiple factors may result in systematically unreliable, unstable AU4 estimation from a person by our AU model. We exclude 3 participants with unreliable AU4 estimation from the analysis of automatic preference annotation with the following criterion: a participant is excluded if the 80th percentile of the AU4 activation values in their baseline clips exceeds 0.5. (Actually, the lowest among the excluded ones is 0.74 while the highest among the included ones is 0.36.) Among the three excluded participants, two of them have bangs covering the area between the eyebrows while the other one has highly unstable camera angles in the videos.
5.2.2 AU4 Valence Score
The AU4 activation value $\alpha_{4}$ is a good indicator of the user’s evaluation of the generated image, but the raw value might not perfectly reflect the associated evaluation. Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image. We define the AU4 valence score $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ as follows:
$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}}),$ (3)
where $k$ is an exponential decay coefficient to fit.
5.2.3 Baseline Scores
To compare the performance of our AU4 valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [40], ImageReward score [73], PickScore [32], and HPS v2 score [71] as baselines.
Among these baseline scoring models, ImageReward, PickScore, and HPS v2 were all trained specifically on datasets of human preferences of text-to-image generation and are supposed to outperform the other models [73, 32, 71]. To introduce a potentially more competitive baseline score, we further combine them to form an ensemble baseline score
$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$
subject $\text{subject to}\quad w_{\mathrm{IR}}+w_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}$ $=1,$ (4)
where $s_{\mathrm{ens}}$ represents the ensemble baseline score, $s_{\mathrm{IR}}$, $s_{\mathrm{Pick}}$, and $s_{\mathrm{HPSv2}}$ represent the ImageReward score, PickScore, and HPS v2 score respectively, and $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$ represent their respective weights.
To be consistent with the evaluation of AU4 valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction. Since the scales of the scores are different, all of them were standardized based on the mean and standard deviation of the training set. We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
5.2.4 AU4 Valence Score: Independent Preference Prediction for Selected Image Pairs
We first evaluate AU4 valence score on its ability to independently predict user preferences between pairs of images generated based on the same text prompt. Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data of a total of 7068 image pairs from the 30 included participants.
Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AU4 valence scores may not be informative for inferring the preference between the two associated images. Therefore, we apply a difference threshold $d$ for making predictions: only when the difference between the AU4 valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AU4 valence scores; otherwise, we abstain from making a prediction.
There are 2 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eq. 3) and the difference threshold $d$. We fit these two parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$ and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
$n_{\mathrm{correct}}-2n_{\mathrm{incorrect}},$ (5)
where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the numbers of correct predictions and incorrect predictions on preference of image pairs respectively.
Model Acc.
CLIP Score 60.58
Aesthetic Score 50.99
BLIP Score 52.98
ImageReward Score 63.11
PickScore 62.57
HPS V2 Score 62.93
Ensemble Baseline Score 67.09
AU4 Valence Score 74.86
Table 1: Accuracy of image preference prediction on the 15.6% of all data (1106 out of 7068 pairs) as selected by the AU4 valence model. (Image pairs with AU4 valence score difference below the trained difference threshold are not included.)
The results are shown in Tab. 1. Based on the trained difference thresholds, the AU4 valence model makes predictions on 15.6% of the image preference data (1106 out of 7068 pairs). For the selected image pairs, our AU4 valence score has the best accuracy of 74.86%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 7.77%.
Figure 4: Heatmap of the image preference prediction accuracy of AU4 valence score on the selected subset of all data. The y-axis represents different values of the exponential decay coefficient in the computation of AU4 valence score, and the x-axis represent different values of the difference threshold. As indicated by the color bars, for the color of each pixel, the hue represents the AU4 valence score prediction accuracy associated with the set of two parameters (exponential decay coefficient and difference threshold) while the saturation represents the proportion of selected data under the set of two parameters. The orange square marks the pixel representing the set of coefficients with optimized Eq. 5 (hue of this pixel is used in the color bar for proportion of selected data). The color bar for accuracy ranges only from 0.5 to 1.0 in our visualization, as no accuracies below 0.5 were obtained.
Our results have high stability. Firstly, the parameter values fit in the LOPO procedure were generally consistent. We obtained the same values of $k=0.4$ and $d=0.42$ for 28 out of the 30 total training sets; for the other two sets, we had $k=0.5$ and $k=0.46$. Secondly, Fig. 4 shows the accuracies of AU4 valence score under different sets of the two parameters (exponential decay coefficient and difference threshold), from which we can clearly see that the accuracy is smooth in these parameters and that high accuracy holds for a wide range of parameters.
5.2.5 Integrate AU4 Valence Score with Pre-trained Scoring Models
AU4 valence score has a great performance of 74.86% on the subset of the image pairs with distinguishable AU4 valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 56.62%, lower than that of the CLIP score, ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab. 2).
However, besides making independent predictions, the AU4 valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data. We integrate the pre-trained scoring models with AU4 valence score with a simple linear combination:
$s_{m}+a_{m}s_{\mathrm{AU4}},$ (6)
where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score given by the model $m$, and $a_{m}$ the weight for integrating model $m$ and the AU4 valence score.
We separately investigated the effect of integrating the AU4 valence score with the three scoring models trained on large human preference datasets, ImageReward, PickScore, and HPS v2, and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 3.0 with a granularity of 0.1) to optimize the accuracy of image preference prediction; for the weights of the ensemble baseline score (Eq. 4) and the exponential decay coefficient for the AU4 valence score (Eq. 3), we directly used the values fit with a similar LOPO procedure in Secs. 5.2.3 and 5.2.4.
Model Ind. Acc. Acc. w/ AU4
CLIP Score 57.50 /
Aesthetic Score 49.92 /
BLIP Score 49.92 /
AU4 Valence Score 56.62 /
ImageReward Score 60.44 62.68
PickScore 63.36 65.31
HPS V2 Score 61.94 65.45
Ensemble Baseline Score 65.79 66.81
Table 2: Accuracy of image preference prediction on all data. The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AU4 valence score.
The results are shown in Tab. 2. Integration with AU4 valence score improves the performance of all of the four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the AU4 valence score, we achieve the highest accuracy of 66.81%, outperforming the ensemble of the three models by a margin of 1.02%.
5.3 Aspects of Evaluations Reflected
We have just compared the performances of different scoring models. An interesting question is what aspects of the evaluation these models respectively reflect. Evaluations on image-text alignment and fidelity (image quality) are the two most essential aspects of text-to-image generation assessment researchers have been focusing on [52]. In our annotation survey, participants give a rating for each of them on a scale of 1 to 7 in addition to the overall rating of the image. To investigate their respective effects on the AU4 valence score, we apply a multiple linear regression from image-text alignment ratings and fidelity ratings to standardized scores from different models across all of the 30 participants included in Sec. 5.2. (AU4 valence scores are computed with a consistent exponential decay coefficient of 0.4; as mentioned in Sec. 5.2.4, the same parameter value of $k=0.4$ was obtained consistently in 28 out of the 30 LOPO training sets.) As shown in Tab. 3, the AU4 valence score has the most significant coefficient for fidelity ratings while the pre-trained scoring models (ImageReward, PickScore, HPS v2, and their ensemble) have more significant coefficient for image-text alignment ratings. This suggests that the aspect of evaluation AU4 primarily reflects is complementary to that of the pre-trained scoring models and explains why integrating the AU4 valence score with them consistently improves their performance.
Model Coef. for Alignment P-val. for Alignment Coef. for Fidelity P-val. for Fidelity
CLIP Score $-3.15\times 10^{-2}$ $5.77\times 10^{-2}$ $7.47\times 10^{-2}$ $2.33\times 10^{-5}$
Aesthetic Score $-8.71\times 10^{-3}$ $6.01\times 10^{-1}$ $1.85\times 10^{-2}$ $2.97\times 10^{-1}$
BLIP Score $-1.95\times 10^{-2}$ $2.41\times 10^{-1}$ $1.72\times 10^{-2}$ $3.31\times 10^{-1}$
ImageReward Score $2.53\times 10^{-1}$ $3.71\times 10^{-57}$ $-5.30\times 10^{-2}$ $1.19\times 10^{-3}$
PickScore $1.53\times 10^{-1}$ $9.06\times 10^{-22}$ $3.54\times 10^{-2}$ $3.53\times 10^{-2}$
HPS V2 Score $6.98\times 10^{-2}$ $2.31\times 10^{-5}$ $3.49\times 10^{-2}$ $4.64\times 10^{-2}$
Ensemble Baseline Score $1.38\times 10^{-1}$ $6.64\times 10^{-24}$ $2.64\times 10^{-2}$ $6.67\times 10^{-2}$
AU4 Valence Score $3.57\times 10^{-2}$ $2.89\times 10^{-2}$ $7.50\times 10^{-2}$ $1.64\times 10^{-5}$
Table 3: The results of the multiple linear regression from the image-text alignment ratings and fidelity ratings to standardized scores from different models. The five columns represent the name of the model, the coefficient for image-text alignment rating, the p-value for image-text alignment rating, the coefficient for fidelity rating, and the p-value for fidelity rating respectively.
6 Related Work
Text-to-Image Generation and Evaluation. Various models have been developed for text-to-image generation, including Generative Adversarial Networks (GANs) [22, 29, 3, 28, 13, 14], Variational autoencoders (VAEs) [31, 8, 64], flow-based models [15, 16, 30], autoregressive models (ARMs) [7, 9, 67, 68, 78], and diffusion models (DMs) [62, 25, 63]. More recent advancements in DMs [50, 58, 57, 54, 59, 12] have achieved great success in generating high-quality images and attracted widespread public attention. As traditional evaluation metrics like Inception Score (IS) [2], Fréchet Inception Distance (FID) [24], and CLIP score [55] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [72], ImageReward [73], PickScore [32], and Human Preference Score v2 (HPS v2) [71], and demonstrated the effectiveness of using them to improve the text-to-image generative models [73, 32, 71, 21].
Human Feedback Datasets. There are multiple datasets of human feedback on text-to-image generation, including AGIQA-1K [80], Human Preference Dataset (HPD) [72], ImageReward [73], Pick-a-Pic [32], AGIQA-3K [39], and Human Preference Dataset v2 (HPD v2) [71]. To the best of our knowledge, our FERGI dataset is the first to include both manual annotations of generated images and the associated facial expression reaction videos.
Facial Expression Recognition. Automatic facial expression recognition has advanced rapidly in recent years [41]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [18, 17, 46], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [19]. AU detection and estimation has attracted increasing interest [26, 61, 82, 81, 10, 42, 43, 69, 33, 34, 83, 45, 65, 66] and been used for facial emotion recognition [75, 77, 76, 74]. Specifically related to our usage of AU4 to infer user evaluations of generated images, researchers have shown that the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [23] and increases with negative stimuli, such as negative images [5, 4, 37, 36].
7 Discussion and Conclusion
We propose that annotation of user preferences for text-to-image generation can be automated with analysis of user facial expression reaction to generated images, which can potentially increase the efficiency and scalability of human preference data collection. We present the FERGI dataset, comprising of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
Specifically, AU4 activation is most consistently reflective of negative evaluation of the image, so we developed an AU4 valence score to represent the user’s evaluation of the image reflected by AU4 activation, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt. Firstly, the AU4 valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including the ensemble of them. Therefore, it can potentially be used to selectively annotate large-scale user preference data to be used for fine-tuning scoring models that do not rely on data of facial expression reaction. Secondly, the AU4 valence score can be integrated with pre-trained scoring models to improve their overall accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label data of large scale human preferences for fine-tuning text-to-image generative models. Thirdly, AU4 may be useful for understanding individual user preferences to create personalized image generation models. Additionally, we show that the AU4 valence score is the most reflective of the evaluation of the fidelity of the generated images among all models, which is complementary to the pre-trained scoring models, which better reflect the evaluation of image-text alignment.
Finally, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction in this work. The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [53], image inpainting [20], and super-resolution [70].
8 Acknowledgments
We thank Xiaojing Xu and Yuan Tang for helpful prior work and Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions. We are grateful for support from NSF IIS 1817226 and seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion as well as hardware funding from NVIDIA, Adobe, and Sony.
Supplementary Material
A Supplementary Visuals
Figure S1: A flow chart summarizing the pipeline our paper proposes. The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated in the paper. Specifically, we developed a model and method for computing an AU4 valence score and integrating it with state-of-the-art text-to-image scoring models. We propose that the AU4 valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to improve/fine-tune the text-to-image generative models.
Figure S2: An example of AU4 activation in response to a low-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity). In the annotation survey, the participant gives an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she also ranked the image as the worst among the five images generated from the same prompt.
To better illustrate the idea of our paper, we present Fig. S1 showing a flow chart summarizing the pipeline the paper proposed and Fig. S2 showing an example of a strong AU4 activation in response to a low-quality image generation.
B Data Collection
This section provides more details of the collection of the FERGI dataset not provided in Sec. 2. Figs. S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
B.1 Participants
39 participants [^1] The user IDs are up to 040 as shown in Tab. S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that is available upon request for research purposes.
B.2 Prompt Creation
There are two forms of sessions, structured input sessions and free-form input sessions. Most participants completed the same number of structured input sessions and free-input sessions. The number of sessions completed by each participant are shown in Tab. S1. By default, each participant is expected to complete 9 structured input sessions and 9 free-form input sessions (for the first 13 out of the 33 participants chronologically) or to complete 10 structured input sessions and 10 free-form input sessions (for the last 20 out of the 33 participants chronologically). However, the time consumption highly varies for each participant because of numerous reasons (e.g. Internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
User ID Number of Structured Input Sessions Number of Free-Form Input Sessions Number of Valid Images
001 10 10 100
002 5 5 50
004 9 9 87
005 9 9 85
006 9 9 90
007 9 9 88
008 9 9 85
009 9 9 88
010 9 8 84
011 9 9 87
012 8 8 80
015 9 9 89
016 10 10 100
017 3 0 15
018 11 4 75
020 10 10 92
021 9 9 87
022 9 6 75
023 10 10 99
024 10 2 58
025 3 10 63
026 10 10 100
028 0 20 99
030 10 10 99
031 10 10 100
032 10 10 99
033 10 10 99
034 10 10 98
035 10 10 99
037 10 10 100
038 10 10 96
039 10 10 96
040 9 4 65
Total 288 288 2827
Table S1: Information of the FERGI dataset.
B.2.1 Structured Input
The purpose of having structured input sessions is to encourage creation of more diverse text prompts featuring different types of objects, interactions, settings, and styles. Structured input contains 5 sections:
  • Animate Objects (Fig. 2(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc. Participants are required to enter the name and quantity of the object, with activity and a list of characteristics as optional inputs.
  • Inanimate Objects (Fig. 2(b)): Any object that cannot do things, including plants, vehicles, furniture, etc. Participants are required to enter the name and quantity of the object, with a list of characteristics as optional inputs.
  • Interactional Relations (Fig. 2(c)): Relationship between an animated object and any other object (something an animate object does to another object). An animate object can only be in a relation if its “activity” field is empty. All objects can only be involved in at most one relation (either interactional or positional).
  • Positional Relations (Fig. 2(d)): Relationship between any two objects. An animate object can only be in a relation if its “activity” field is empty. All objects can only be involved in at most one relation (either interactional or positional).
  • Other Inputs (Fig. S4): The location of image as background or general environment, the style of the image, and a list of keywords to append at the end of the prompt.
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs. S3 and S4.
B.2.2 Free-Form Input
The purpose of free-form input is to provide freedom and flexibility for participants to experiment with any prompts to complement the less flexible input style of structured input. The participants are provided with an example prompt and a text box to enter the entire prompt as the final prompt. A screenshot of an example of the prompt creation page for a free-form input session is shown in Fig. S5.
B.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
B.3.1 Image Annotation
The annotation survey was adapted from the survey designed for ImageReward [73] and has the following components:
  • Participants can optionally provide a typed, comma-separated response indicating any phrase or parts of the prompt that are not reflected by the generated image.
  • Participants provide star ratings from 1 to 7 for image-text alignment, fidelity, and overall rating for the generated image.
  • Participants can optionally select from a list of common issues to report the issues the generated image has.
  • Participants can optionally select from a list of emotions to report their feelings when seeing the image.
A screenshot of an example of the image annotation page (including the details of the questions) is shown in Fig. S8.
B.3.2 Image Ranking
After the participants finish annotating all 5 generated images, they will be asked to rank them from best to worst. All 5 images are displayed side by side, with the left side labeled “Best” and right side labeled “Worst”, and the participants provide ranking by dragging the images to reorder them. A screenshot of an example of the image ranking page is shown in Fig. S9.
(a) Animate objects
(b) Inanimate objects
(c) Interactional relations
(d) Positional relations
Figure S3: An example of the basic inputs for the prompt creation of structured input sessions.
Figure S4: An example of other inputs and prompt generation for the prompt creation of structured input sessions.
Figure S5: An example of the prompt creation page for the free-form input sessions.
Figure S6: An example of a baseline page.
Figure S7: An example of an image presentation page.
Figure S8: An example of an image annotation survey page.
Figure S9: An example of an image ranking page.
C AU Model Training
This section provides more details of the training of the AU model not provided in Sec. 3.
C.1 Datasets
The DISFA dataset [49] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total. Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5222See https://melindaozel.com/facs-cheat-sheet/ for a visual reference guide for FACS.
The DISFA+ dataset [47] is an extension of the DISFA dataset [49]. It contains facial video recordings of 9 participants’ posed and spontaneous facial expression with each frame being annotated with the same 12 AUs on a scale of 0 to 5.
C.2 Video Preprocessing
For each frame of the video, we first detect and crop the largest face in the image using MediaPipe [44] and then align the face with Pytorch Face Landmark Detection [6]. Then, we further employ a combination of histogram equalization and linear mapping [35] to increase the global contrast of the facial image. The facial images are resized to $112\times 112$ pixels before being fed into the neural network. Similar preprocessing is also applied to the videos we analyze in the FERGI dataset.
C.3 Model Training
We use the neural network IR-50 [11] pre-trained on Glink360k [1] and fine-tune it on the DISFA and DISFA+ datasets [49, 47]. The last layer of the network is modified so that it outputs the estimation of the AUs in two formats: for estimating the intensity of the $i$th AU $y_{i}$, it outputs 1 value $\hat{y}_{i,\mathrm{reg}}$ representing the numerical estimation of the AU intensity (in the format of regression) and 5 values $\hat{y}_{i,\mathrm{class(1)}}$, $\hat{y}_{i,\mathrm{class(2)}}$, $\hat{y}_{i,\mathrm{class(3)}}$, $\hat{y}_{i,\mathrm{class(4)}}$, and $\hat{y}_{i,\mathrm{class(5)}}$ respectively representing the estimated probability of the AU intensity being higher than or equal to 1, 2, 3, 4, and 5 [51] (in the format of binary classifications). The loss function consists of three parts:
$E=E_{\mathrm{reg,MSE}}+E_{\mathrm{reg,cos}}+E_{\mathrm{class}},$ (7)
where $E_{\mathrm{reg,MSE}}$, $E_{\mathrm{reg,cos}}$, and $E_{\mathrm{class}}$ respectively represent a mean squared error (MSE) loss for the numerical estimations
$E_{\mathrm{reg,MSE}}=\Sigma_{i=1}^{n}w_{i,y_{i}}(y_{i}-\hat{y}_{i,\mathrm{reg}})^{2},$ (8)
a cosine similarity loss for the numerical estimations
$E_{\mathrm{reg,cos}}=1-\frac{\Sigma_{i=1}^{n}y_{i}\hat{y}_{i,\mathrm{reg}}}{(\Sigma_{i=1}^{n}{y_{i}^{2}})(\Sigma_{i=1}^{n}\hat{y}_{i,\mathrm{reg}}^{2})},$ (9)
and a cross entropy loss for the binary classification estimations
$E_{\mathrm{class}}=\Sigma_{i=1}^{n}\Sigma_{j=1}^{5}w_{i,j,\chi_{y_{i}\geq j}}CE(\chi_{y_{i}\geq j},\sigma(\hat{y}_{i,\mathrm{class(j)}})),$ (10)
with the cross entropy function being
$CE(y,\hat{y})=-[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})].$ (11)
The weights for the MSE loss and those for the cross entropy loss are both inverse-frequency weighted (for the MSE loss these are based on the merged groups of $\{0,1\}$ and $\{2,3,4,5\}$). Specifically, the weights for the MSE loss are defined as
$w_{i,j}=\begin{cases}\frac{2\cdot\frac{1}{\Sigma_{j^{\prime}=0}^{1}n_{i,j^{\prime}}}}{2\cdot\frac{1}{\Sigma_{j^{\prime}=0}^{1}n_{i,j^{\prime}}}+4\cdot\frac{1}{\Sigma_{j^{\prime}=2}^{5}n_{i,j^{\prime}}}},&\text{for $j=0,1$}\\ \frac{4\cdot\frac{1}{\Sigma_{j^{\prime}=2}^{5}n_{i,j^{\prime}}}}{2\cdot\frac{1}{\Sigma_{j^{\prime}=0}^{1}n_{i,j^{\prime}}}+4\cdot\frac{1}{\Sigma_{j^{\prime}=2}^{5}n_{i,j^{\prime}}}},&\text{for $j=2,3,4,5$}\end{cases}$ (12)
while the weights for the cross entropy loss are defined as
$\begin{cases}w_{i,j,1}=\frac{\frac{1}{\Sigma_{j^{\prime}=j}^{5}n_{i,j^{\prime}}}}{\Sigma_{j^{\prime\prime}=1}^{5}(\frac{1}{\Sigma_{j^{\prime}=0}^{j^{\prime\prime}-1}n_{i,j^{\prime}}}+\frac{1}{\Sigma_{j^{\prime}=j^{\prime\prime}}^{5}n_{i,j^{\prime}}})}\\ w_{i,j,0}=\frac{\frac{1}{\Sigma_{j^{\prime}=0}^{j-1}n_{i,j^{\prime}}}}{\Sigma_{j^{\prime\prime}=1}^{5}(\frac{1}{\Sigma_{j^{\prime}=0}^{j^{\prime\prime}-1}n_{i,j^{\prime}}}+\frac{1}{\Sigma_{j^{\prime}=j^{\prime\prime}}^{5}n_{i,j^{\prime}}})},\end{cases}$ (13)
where $n_{i,j}$ represents the number occurrences of the $i$th AU with an intensity of $j$.
Notably, although we train the neural network to learn both numerical estimations and binary classification estimations of the AU intensities, only the numerical estimations are used in model inference.
For model training, we employ the Adam optimizer with an initial learning rate of $10^{-4}$ for parameters of the last layer and an $10^{-5}$ for other parameters, a weight decay of $5\times 10^{-4}$, and a batch size of 64. We train the model on all data from the DISFA and DISFA+ datasets [49, 47] for a total of 3 epochs using a single NVIDIA GeForce GTX 1080Ti 11G GPU.
D Data Filtering
This section provides the details of frame exclusion not provided in Sec. 4.1.
We aim to exclude frames with occlusions and frames with off-angle pitch or yaw. To achieve the goal, we follow the following three exclusion rules:
  • Frames with a face detection confidence score (FDCS) lower than 0.9 are excluded. The FDCS is given by MediaPipe [44]. Low FDCS is likely caused by occlusions.
  • Frames with a yaw indicative ratio (YIR) out of the range of $[0.3,0.7]$. The YIR is computed as
    $\textrm{YIR}=\frac{d_{\textrm{eye-edge,left}}}{d_{\textrm{eye-edge,left}}+d_{\textrm{eye-edge,right}}},$ (14)
    where $d_{\textrm{eye-edge,left}}$ represents the horizontal distance between the left eye and the left edge of the face while $d_{\textrm{eye-edge,right}}$ represents the horizontal distance between the right eye and the right edge of the face, both of which are computed based on the facial landmarks detected with [6].
  • Frames with a pitch indicative ratio (PIR) out of the range of $[0.55,0.85]$. The PIR is computed as
    $\textrm{PIR}=\frac{d_{\textrm{nostrils-eyes}}}{d_{\textrm{eyes}}},$ (15)
    where $d_{\textrm{nostrils-eyes}}$ represents the vertical distance between the center of the nostrils and the center of the eyes while $d_{\textrm{eyes}}$ represents the horizontal distance between the two eyes, both of which are computed based on the facial landmarks detected with [6].

### v2

@@
 We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images.
-Specifically, AU4 (brow lowerer) is most consistently reflective of negative evaluations of the generated image.
-This can be useful in two ways.
-Firstly, we can automatically annotate user preferences between image pairs with substantial difference in AU4 responses to them with an accuracy significantly outperforming state-of-the-art scoring models.
-Secondly, directly integrating the AU4 responses with the scoring models improves their consistency with human preferences.
-Additionally, the AU4 response best reflects the user’s evaluation of the image fidelity, making it complementary to the state-of-the-art scoring models, which are generally better at reflecting image-text alignment.
+Specifically, AU4 (brow lowerer) is reflective of negative evaluations of the generated image whereas AU12 (lip corner puller) is reflective of positive evaluations.
+These can be useful in two ways.
+Firstly, we can automatically annotate user preferences between image pairs with substantial difference in these AU responses with an accuracy significantly outperforming state-of-the-art scoring models.
+Secondly, directly integrating the AU responses with the scoring models improves their consistency with human preferences.
 Finally, this method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks.
@@
 This automatic annotation requires zero additional effort from the real users of text-to-image generators and is thus expected to increase the scalability of human preference feedback data collection.
-We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion v1.4 [58] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input.
+We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion (SD) v1.4 [58] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input.
 We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [20]), in the facial expression reaction videos with a pre-trained model and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images.
-We further propose a method of automatically annotating user preferences by detecting the activation of AU4 in the facial expression reaction to the generated images and show that this evaluation metric best reflects the user’s evaluation of the image fidelity, making it complementary to the state-of-the-art scoring models, which are generally better at reflecting image-text alignment.
+We further propose a method of automatically annotating user preferences by detecting the activations of AU4 (brow lowerer, see Fig.
+5) and AU12 (lip corner puller, see Fig.
+6) in the facial expression reaction to the generated images.
+Figure 1: A flow chart summarizing our proposed pipeline.
+The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated.
+Specifically, we developed a model and method for computing an AUcomb valence score and integrating it with state-of-the-art text-to-image scoring models.
+We propose that the AUcomb valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to fine-tune the text-to-image generative models.
+To better illustrate the idea of our paper, we present Fig. 1 showing a flow chart summarizing the pipeline the paper proposed.
+2 Related Work
+Text-to-Image Generation and Evaluation.
+Various models have been developed for text-to-image generation, including Generative Adversarial Networks (GANs) [23, 30, 4, 29, 14, 15], Variational autoencoders (VAEs) [32, 9, 64], flow-based models [16, 17, 31], autoregressive models (ARMs) [8, 10, 67, 68, 80], and diffusion models (DMs) [62, 26, 63].
+More recent advancements in DMs [51, 58, 57, 54, 59, 13] have achieved great success in generating high-quality images and attracted widespread public attention.
+As traditional evaluation metrics like Inception Score (IS) [2], Fréchet Inception Distance (FID) [25], and CLIP score [55] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [74], ImageReward [75], PickScore [33], and Human Preference Score v2 (HPS v2) [73], and demonstrated the effectiveness of using them to improve the text-to-image generative models [75, 33, 73, 22].
+Human Feedback Datasets.
+There are multiple datasets of human feedback on text-to-image generation, including AGIQA-1K [82], Human Preference Dataset (HPD) [74], ImageReward [75], Pick-a-Pic [33], AGIQA-3K [40], and Human Preference Dataset v2 (HPD v2) [73].
+To the best of our knowledge, our FERGI dataset is the first to include both manual annotations of generated images and the associated facial expression reaction videos.
+Facial Expression Recognition and Applications.
+Automatic facial expression recognition has advanced rapidly in recent years [42].
+As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [19, 18, 47], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [20].
+AU detection and estimation has attracted increasing interest [27, 61, 84, 83, 11, 43, 44, 69, 34, 35, 85, 46, 65, 66] and been used for facial emotion recognition [77, 79, 78, 76].
+Specifically related to our usage of AU4 and AU12 to infer user evaluations of generated images, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [72].
+Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [24] and increases with negative stimuli, such as negative images [6, 5, 38, 37].
+For applications, [3, 71] showed the effectiveness of using facial expression to infer/analyze human preferences.
 3 FERGI Dataset
@@
 3.2 Data Collection Procedure
-Each participant completes multiple sessions of data collection.
-In each session, the participant drafts one text prompt and views 5 images generated from the input text prompt.
+Each participant completed multiple sessions of data collection.
+In each session, the participant drafted one text prompt and viewed 5 images generated from the input text prompt.
 A flow chart of the procedure of each session is shown in Fig.
-1.
+2.
 Details are explained below.
@@
 The webcam is then used to record throughout the rest of the session.
-The configuration of the raw videos vary based on the participant’s device and browser.
-Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos are standardized to 30 FPS afterwards.
+The configuration of the raw videos varied based on the participant’s device and browser.
+Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos were standardized to 30 FPS afterwards.
 3.2.3 Image Presentation, Annotation, and Ranking
-In each session, a total of 5 images generated by Stable Diffusion v1.4 [58] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images).
-The image presentation and annotation follows the webcam preview.
+In each session, a total of 5 images generated by SD v1.4 [58] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images).
+The image presentation and annotation followed the webcam preview.
 The detailed procedures for the presentation and annotation of each generated image are as follows:
@@
 The participant annotates the image manually by filling out the survey during this interval.
-The survey was adapted from the survey designed for ImageReward [73] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (which was fixed afterwards) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis.
+The survey was adapted from the survey designed for ImageReward [75] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (later fixed) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis.
 (multi-choice question), and emotions felt when seeing the image (multi-choice question).
-After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server.
+After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server. That marks the end of the session.
 3.3 Data Preprocessing
 For each image, two clips are extracted from the recorded video: a 5-second “baseline clip” of the participant’s facial expression during the baseline period and a 5-second “reaction clip” of the participant’s facial expression during the image presentation period.
-Stable Diffusion v1.4 [58] outputs an all-black image when it detects the original output image might be inappropriate.
+SD v1.4 [58] outputs an all-black image when it detects the original output image might be inappropriate.
 A total of 53 such images out of 2880 total generated were generated in the dataset.
-The data associated with these images were excluded from further analysis, resulting in less than 5 valid images for some input text prompts.
+The data associated with these images were excluded from further analysis, resulting in fewer than 5 valid images for some input text prompts.
 Figure 2: Procedure of one session in data collection.
@@
 We trained our AU estimation model on the DISFA [50, 49] and DISFA+ [48] datasets.
+They are among the largest, most popular, and most cited AU datasets with intensity annotations.
 In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop).
 Video frames are first preprocessed with face detection [45], face alignment [7], and a combination of histogram equalization and linear mapping [36] before being fed into the model for training, testing, or inference.
-For model training, we use the neural network IR-50 [11] pre-trained on Glink360k [1] and fine-tune it on the DISFA and DISFA+ datasets [49, 47].
-A single network learns the estimation of all AUs labeled in the datasets with both regression and ordinal classification [51].
+For model training, we used the neural network IR-50 [12] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [50, 48].
+A single network learned the estimation of all labeled AUs with both regression and ordinal classification [52].
+More details of the training of the AU model is elaborated in the Supplementary Materials.
 5 Facial Feature Extraction
@@
 Similar procedures were used to extract features from baseline clips and reaction clips.
+Further details are in the Supplemental Materials.
 5.1 Data Filtering
 Robustness against face occlusion and faces with non-frontal poses is an ongoing challenge for facial expression recognition models [81, 28].
-Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and has not been specifically adapted for non-frontal faces or faces with occlusion, so it is not suitable for estimating AUs of such faces.
-We address this issue by excluding such data from analysis.
-Specifically, for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [44]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [6]).
+Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and thus for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [45]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [7]).
 Then, if a clip has more than $20\%$ of its total frames being excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
 5.2 AU Activation Value
-We compute an AU activation value $\alpha_{i}$ for each trained AU (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
+We compute an AU activation value $\alpha_{i}$ for each trained AUi (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
   • Firstly, we estimate the intensity of each AU for each frame of the video clip with the AU model: $\hat{y}_{i,\mathrm{reg}}^{(1)},\hat{y}_{i,\mathrm{reg}}^{(2)},\dots$, where the superscript indicates the index of the frame.
@@
 (c) AU1
-(d) AU2
-(e) AU5
-(f) AU12
-(g) AU20
+(d) AU5
+(e) AU20
+(f) AU2
+(g) AU12
 […]
@@
   • Overall Ratings.
-The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs.
-1(a) and 1(b) (Spearman correlation: $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9).
+The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant positive correlation with the activation values of AU2 and AU12 and a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs.
+2(f), 2(a), 2(b) and 2(g) (Spearman correlation: $\rho\approx 0.09$ and $p<5\times 10^{-5}$ for AU2, $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9, $\rho\approx 0.08$ and $p<5\times 10^{-4}$ for AU12).
 […]
   • Emotions.
-For the question “Did you feel any of the following emotions when you saw the image?”, the participant can decide whether to include the six options (disappointed, satisfied, surprised, disgusted, amused, scared) in the answer independently (they can select 0 to 6 of the options as their answer).
-We found that whether they reportedly feel each of the emotions are significantly correlated with the activation values of multiple AUs.
+For the question “Did you feel any of the following emotions when you saw the image?”, the participant can independently select 0 to 6 answers from the six options: disappointed, satisfied, surprised, disgusted, amused, and scared.
+Each of the reported emotions was found to be significantly correlated with the activation values of multiple AUs.
     • Disappointed.
-”Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig.
-2(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
+“Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig.
+3(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
     • Satisfied.
-”Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig.
-2(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
+“Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig.
+3(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
     • Surprised.
-”Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig.
-2(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
+“Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig.
+3(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
     • Disgusted.
-”Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig.
-2(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
+“Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig.
+3(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
     • Amused.
-”Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig.
-2(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
+“Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig.
+3(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
     • Scared.
-”Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig.
-2(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
+“Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig.
+3(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
 Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$, with 12 representing the 12 investigated AUs, 2 representing the two tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 representing the 6 tests for 6 emotions.
-5.2 Automatic Annotation of User Preferences between Image Pairs Using AU4
-We have observed that AU4 activation value has the highest and most consistent (negative) correlation with the raw overall ratings in the previous subsection.
-Therefore, we use the participants’ AU4 activation values to automatically annotate their preferences between images generated based on the same input text prompt.
+6.2 Automatic Annotation of User Preferences between Image Pairs Using AU4 and AU12
+Figure 5: An example of AU4 activation in response to a low-quality image generation.
+(a) The input text prompt.
+(b) The facial expression of the participant before seeing the generated image.
+(c) The generated image.
+(d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity).
+In the annotation survey, the participant gave an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she ranked the image as the worst among the five images generated from the same prompt.
+Figure 6: An example of AU12 activation in response to a high-quality image generation.
+(a) The input text prompt.
+(b) The facial expression of the participant before seeing the generated image.
+(c) The generated image.
+(d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU12 intensity).
+In the annotation survey, the participant gave ratings of 7 for each of image-text alignment, fidelity, and overall rating; he did not identify any issues with the image; he reported feeling “satisfied”, “surprised”, and “amused” when seeing the image; he ranked the image as the best among the five images generated from the same prompt.
+We aim to use the estimated AU activation values to automatically annotate user preferences between images generated from the same input text prompt.
+We have observed that the activations of AU2, AU4, AU9, and AU12 are significantly correlated with the the raw overall ratings in the previous subsection.
+After comparing combinations between them, we developed the following pipeline of using AU4 and AU12 to do automatic annotation (examples of AU4 and AU12 responses are shown in Figs.
+5 and 6).
+(Inclusion of AU2 and AU9 did not help.) We also did not find a way to effectively incorporate the information conveyed by AUs that correlate only with the extremity of the ratings.
+Collection of a much larger-scale dataset in the future might enable effective incorporation of such information with a more complex model.
 6.2.1 Participant Exclusion
-Multiple factors may result in systematically unreliable, unstable AU4 estimation from a person by our AU model.
-We exclude 3 participants with unreliable AU4 estimation from the analysis of automatic preference annotation with the following criterion: a participant is excluded if the 80th percentile of the AU4 activation values in their baseline clips exceeds 0.5.
-(Actually, the lowest among the excluded ones is 0.74 while the highest among the included ones is 0.36.) Among the three excluded participants, two of them have bangs covering the area between the eyebrows while the other one has highly unstable camera angles in the videos.
-5.2.2 AU4 Valence Score
-The AU4 activation value $\alpha_{4}$ is a good indicator of the user’s evaluation of the generated image, but the raw value might not perfectly reflect the associated evaluation.
-Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image.
-We define the AU4 valence score $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ as follows:
-$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}}),$ (3)
-where $k$ is an exponential decay coefficient to fit.
+While estimations of AU12 are generally reliable and stable, multiple factors may result in unreliable AU4 estimation from a small number of people by our AU model.
+We excluded 3 participants with unreliable AU4 estimation due to bangs covering the area between the eyebrows (2 participants) or highly unstable camera angles (1 participant).
+6.2.2 AUcomb Valence Score
+The AU4 activation value $\alpha_{4}$ and AU12 activation value $\alpha_{12}$ are good indicators of the user’s evaluation of the generated image, but the raw values might not linearly reflect the associated evaluation.
+Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$ and an AU12 valence score $s_{\mathrm{AU12}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image.
+We define the $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ and $s_{\mathrm{AU12}}$ as the positive exponential decay function of AU12 activation value $\alpha_{12}$ as follows:
+$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}})$ (3)
+and
+$s_{\mathrm{AU12}}=(1-e^{-k\cdot\alpha_{12}}),$ (4)
+where $k$ is an exponential decay coefficient to fit. We further define the final AUcomb valence score $s_{\mathrm{AUcomb}}$ as the linear combination of them:
+$s_{\mathrm{AUcomb}}=w_{\mathrm{AU4}}s_{\mathrm{AU4}}+w_{\mathrm{AU12}}s_{\mathrm{AU12}}$ (5)
+subject $\text{subject to}\quad w_{\mathrm{AU4}}+w_{\mathrm{AU12}}$ $=1,$ (6)
+where $w_{\mathrm{AU4}}$ and $w_{\mathrm{AU12}}$ represent the weights for AU4 and AU12 respectively.
 6.2.3 Baseline Scores
-To compare the performance of our AU4 valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [40], ImageReward score [73], PickScore [32], and HPS v2 score [71] as baselines.
+To compare the performance of our AU valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [41], ImageReward score [75], PickScore [33], and HPS v2 score [73] as baselines.
 Among these baseline scoring models, ImageReward, PickScore, and HPS v2 were all trained specifically on datasets of human preferences of text-to-image generation and are supposed to outperform the other models [75, 33, 73].
 To introduce a potentially more competitive baseline score, we further combine them to form an ensemble baseline score
-$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$
+$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$ (7)
 subject $\text{subject to}\quad w_{\mathrm{IR}}+w_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}$ $=1,$ (8)
 where $s_{\mathrm{ens}}$ represents the ensemble baseline score, $s_{\mathrm{IR}}$, $s_{\mathrm{Pick}}$, and $s_{\mathrm{HPSv2}}$ represent the ImageReward score, PickScore, and HPS v2 score respectively, and $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$ represent their respective weights.
-To be consistent with the evaluation of AU4 valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction.
-Since the scales of the scores are different, all of them were standardized based on the mean and standard deviation of the training set.
+To be consistent with the evaluation of AUcomb valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction.
+Since the scales of the scores are different, they were each z-scored using their mean and standard deviation on the training set.
 We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
-5.2.4 AU4 Valence Score: Independent Preference Prediction for Selected Image Pairs
-We first evaluate AU4 valence score on its ability to independently predict user preferences between pairs of images generated based on the same text prompt.
-Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data of a total of 7068 image pairs from the 30 included participants.
-Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AU4 valence scores may not be informative for inferring the preference between the two associated images.
-Therefore, we apply a difference threshold $d$ for making predictions: only when the difference between the AU4 valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AU4 valence scores; otherwise, we abstain from making a prediction.
-There are 2 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eq.
-3) and the difference threshold $d$.
-We fit these two parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$ and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
+6.2.4 AUcomb Valence Score: Independent Preference Prediction for Selected Image Pairs
+We first evaluate the AUcomb valence score on its ability to independently predict user preferences between pairs of images generated by the same text prompt.
+Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data on a total of 3534 image pairs from the 30 included participants.
+Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AUcomb valence scores may not be informative for inferring preference between the two associated images.
+Therefore, we apply a threshold $d$ for making predictions: only when the difference between the AUcomb valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AUcomb valence score; otherwise, we abstain from making a prediction.
+There are 3 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eqs.
+3 and 4), the weight for AU4 $w_{\mathrm{AU4}}$ (as in Eq.
+5; $w_{\mathrm{AU12}}$ is then automatically set by Eq.
+6) and the difference threshold $d$.
+We fit these three parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$, from 0.0 to 1.0 with a granularity of 0.1 for $w_{\mathrm{AU4}}$, and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
 $n_{\mathrm{correct}}-2n_{\mathrm{incorrect}},$ (9)
-where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the numbers of correct predictions and incorrect predictions on preference of image pairs respectively.
+where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the number of correct predictions and incorrect predictions on image pair preference.
+Table 1: Accuracy of image preference prediction on the 19.38% of all data (685 out of 3534 pairs) as selected by the AUcomb valence model.
+(Image pairs with AUcomb valence score difference below the trained difference threshold are not included.)
 Model Acc.
-CLIP Score 60.58
-Aesthetic Score 50.99
-BLIP Score 52.98
-ImageReward Score 63.11
-PickScore 62.57
-HPS V2 Score 62.93
-Ensemble Baseline Score 67.09
-AU4 Valence Score 74.86
-Table 1: Accuracy of image preference prediction on the 15.6% of all data (1106 out of 7068 pairs) as selected by the AU4 valence model.
-(Image pairs with AU4 valence score difference below the trained difference threshold are not included.)
+CLIP Score 58.25
+Aesthetic Score 51.39
+BLIP Score 49.78
+ImageReward Score 62.19
+PickScore 61.31
+HPS V2 Score 62.04
+Ensemble Baseline Score 66.57
+AUcomb Valence Score 72.70
 The results are shown in Tab.
 1.
-Based on the trained difference thresholds, the AU4 valence model makes predictions on 15.6% of the image preference data (1106 out of 7068 pairs).
-For the selected image pairs, our AU4 valence score has the best accuracy of 74.86%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 7.77%.
-Figure 4: Heatmap of the image preference prediction accuracy of AU4 valence score on the selected subset of all data.
-The y-axis represents different values of the exponential decay coefficient in the computation of AU4 valence score, and the x-axis represent different values of the difference threshold.
-As indicated by the color bars, for the color of each pixel, the hue represents the AU4 valence score prediction accuracy associated with the set of two parameters (exponential decay coefficient and difference threshold) while the saturation represents the proportion of selected data under the set of two parameters.
-The orange square marks the pixel representing the set of coefficients with optimized Eq.
-5 (hue of this pixel is used in the color bar for proportion of selected data).
-The color bar for accuracy ranges only from 0.5 to 1.0 in our visualization, as no accuracies below 0.5 were obtained.
+Based on the trained difference thresholds, the AUcomb valence model makes predictions on 19.38% of the image preference data (685 out of 3534 pairs).
+For the selected image pairs, our AUcomb valence score has the best accuracy of 72.70%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 6.13%.
 Our results have high stability.
-Firstly, the parameter values fit in the LOPO procedure were generally consistent.
-We obtained the same values of $k=0.4$ and $d=0.42$ for 28 out of the 30 total training sets; for the other two sets, we had $k=0.5$ and $k=0.46$.
-Secondly, Fig.
-4 shows the accuracies of AU4 valence score under different sets of the two parameters (exponential decay coefficient and difference threshold), from which we can clearly see that the accuracy is smooth in these parameters and that high accuracy holds for a wide range of parameters.
-5.2.5 Integrate AU4 Valence Score with Pre-trained Scoring Models
-AU4 valence score has a great performance of 74.86% on the subset of the image pairs with distinguishable AU4 valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 56.62%, lower than that of the CLIP score, ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab.
+The parameter values fit in the LOPO procedure were generally consistent.
+We obtained $k=0.5$, $w_{\mathrm{AU4}}=0.5$, $w_{\mathrm{AU12}}=0.5$, and $d=0.26$ for 20 out of the 30 total training sets.
+6.2.5 Integrate AUcomb Valence Score with Pre-trained Scoring Models
+AUcomb valence score has a great performance of 72.70% on the subset of the image pairs with distinguishable AUcomb valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 58.23%, still significantly better than chance (two-sided binomial test: $p<5\times 10^{-22}$) but lower than that of the ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab.
 2).
-However, besides making independent predictions, the AU4 valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data.
-We integrate the pre-trained scoring models with AU4 valence score with a simple linear combination:
-$s_{m}+a_{m}s_{\mathrm{AU4}},$ (6)
-where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score given by the model $m$, and $a_{m}$ the weight for integrating model $m$ and the AU4 valence score.
-We separately investigated the effect of integrating the AU4 valence score with the three scoring models trained on large human preference datasets, ImageReward, PickScore, and HPS v2, and their ensemble.
-Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 3.0 with a granularity of 0.1) to optimize the accuracy of image preference prediction; for the weights of the ensemble baseline score (Eq.
-4) and the exponential decay coefficient for the AU4 valence score (Eq.
-3), we directly used the values fit with a similar LOPO procedure in Secs.
-5.2.3 and 5.2.4.
-Model Ind. Acc. Acc. w/ AU4
+However, besides making independent predictions, the AUcomb valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data.
+We integrate the pre-trained scoring models with AUcomb valence score with a simple linear combination:
+$s_{m}+a_{m}s_{\mathrm{AUcomb}},$ (10)
+where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score from model $m$, and $a_{m}$ the weight for integrating $m$ and the AUcomb valence score.
+We separately investigated the effect of integrating the AUcomb valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble.
+Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); for the weights of the ensemble baseline score (Eq.
+7), exponential decay coefficient for the AU4 and AU12 valence scores (Eqs.
+3 and 4), and the weighting between AU4 and AU12, we directly used the values fit previously.
+Table 2: Accuracy of image preference prediction on all data.
+The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AUcomb valence score.
+Model Ind. Acc. Acc. w/ AUs
 CLIP Score 57.50 /
@@
 BLIP Score 49.92 /
-AU4 Valence Score 56.62 /
-ImageReward Score 60.44 62.68
-PickScore 63.36 65.31
-HPS V2 Score 61.94 65.45
-Ensemble Baseline Score 65.79 66.81
-Table 2: Accuracy of image preference prediction on all data.
-The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AU4 valence score.
+AUcomb Valence Score 58.23 /
+ImageReward Score 60.44 63.67
+PickScore 63.36 65.42
+HPS V2 Score 61.94 65.53
+Ensemble Baseline Score 65.79 67.49
 The results are shown in Tab.
 2.
-Integration with AU4 valence score improves the performance of all of the four baseline models.
-Specifically, by further integrating all three pre-trained human preference scoring models together with the AU4 valence score, we achieve the highest accuracy of 66.81%, outperforming the ensemble of the three models by a margin of 1.02%.
-5.3 Aspects of Evaluations Reflected
-We have just compared the performances of different scoring models.
-An interesting question is what aspects of the evaluation these models respectively reflect.
-Evaluations on image-text alignment and fidelity (image quality) are the two most essential aspects of text-to-image generation assessment researchers have been focusing on [52].
-In our annotation survey, participants give a rating for each of them on a scale of 1 to 7 in addition to the overall rating of the image.
-To investigate their respective effects on the AU4 valence score, we apply a multiple linear regression from image-text alignment ratings and fidelity ratings to standardized scores from different models across all of the 30 participants included in Sec.
-5.2.
-(AU4 valence scores are computed with a consistent exponential decay coefficient of 0.4; as mentioned in Sec.
-5.2.4, the same parameter value of $k=0.4$ was obtained consistently in 28 out of the 30 LOPO training sets.) As shown in Tab.
-3, the AU4 valence score has the most significant coefficient for fidelity ratings while the pre-trained scoring models (ImageReward, PickScore, HPS v2, and their ensemble) have more significant coefficient for image-text alignment ratings.
-This suggests that the aspect of evaluation AU4 primarily reflects is complementary to that of the pre-trained scoring models and explains why integrating the AU4 valence score with them consistently improves their performance.
-Model Coef. for Alignment P-val. for Alignment Coef. for Fidelity P-val. for Fidelity
-CLIP Score $-3.15\times 10^{-2}$ $5.77\times 10^{-2}$ $7.47\times 10^{-2}$ $2.33\times 10^{-5}$
-Aesthetic Score $-8.71\times 10^{-3}$ $6.01\times 10^{-1}$ $1.85\times 10^{-2}$ $2.97\times 10^{-1}$
-BLIP Score $-1.95\times 10^{-2}$ $2.41\times 10^{-1}$ $1.72\times 10^{-2}$ $3.31\times 10^{-1}$
-ImageReward Score $2.53\times 10^{-1}$ $3.71\times 10^{-57}$ $-5.30\times 10^{-2}$ $1.19\times 10^{-3}$
-PickScore $1.53\times 10^{-1}$ $9.06\times 10^{-22}$ $3.54\times 10^{-2}$ $3.53\times 10^{-2}$
-HPS V2 Score $6.98\times 10^{-2}$ $2.31\times 10^{-5}$ $3.49\times 10^{-2}$ $4.64\times 10^{-2}$
-Ensemble Baseline Score $1.38\times 10^{-1}$ $6.64\times 10^{-24}$ $2.64\times 10^{-2}$ $6.67\times 10^{-2}$
-AU4 Valence Score $3.57\times 10^{-2}$ $2.89\times 10^{-2}$ $7.50\times 10^{-2}$ $1.64\times 10^{-5}$
-Table 3: The results of the multiple linear regression from the image-text alignment ratings and fidelity ratings to standardized scores from different models.
-The five columns represent the name of the model, the coefficient for image-text alignment rating, the p-value for image-text alignment rating, the coefficient for fidelity rating, and the p-value for fidelity rating respectively.
-6 Related Work
-Text-to-Image Generation and Evaluation.
-Various models have been developed for text-to-image generation, including Generative Adversarial Networks (GANs) [22, 29, 3, 28, 13, 14], Variational autoencoders (VAEs) [31, 8, 64], flow-based models [15, 16, 30], autoregressive models (ARMs) [7, 9, 67, 68, 78], and diffusion models (DMs) [62, 25, 63].
-More recent advancements in DMs [50, 58, 57, 54, 59, 12] have achieved great success in generating high-quality images and attracted widespread public attention.
-As traditional evaluation metrics like Inception Score (IS) [2], Fréchet Inception Distance (FID) [24], and CLIP score [55] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [72], ImageReward [73], PickScore [32], and Human Preference Score v2 (HPS v2) [71], and demonstrated the effectiveness of using them to improve the text-to-image generative models [73, 32, 71, 21].
-Human Feedback Datasets.
-There are multiple datasets of human feedback on text-to-image generation, including AGIQA-1K [80], Human Preference Dataset (HPD) [72], ImageReward [73], Pick-a-Pic [32], AGIQA-3K [39], and Human Preference Dataset v2 (HPD v2) [71].
-To the best of our knowledge, our FERGI dataset is the first to include both manual annotations of generated images and the associated facial expression reaction videos.
-Facial Expression Recognition.
-Automatic facial expression recognition has advanced rapidly in recent years [41].
-As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [18, 17, 46], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [19].
-AU detection and estimation has attracted increasing interest [26, 61, 82, 81, 10, 42, 43, 69, 33, 34, 83, 45, 65, 66] and been used for facial emotion recognition [75, 77, 76, 74].
-Specifically related to our usage of AU4 to infer user evaluations of generated images, researchers have shown that the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [23] and increases with negative stimuli, such as negative images [5, 4, 37, 36].
+Integration with AUcomb valence score improves the performance of all of the four baseline models.
+Specifically, by further integrating all three pre-trained human preference scoring models together with the AUcomb valence score, we achieve the highest accuracy of 67.49%, outperforming the ensemble of the three models by a margin of 1.70%.
 7 Discussion and Conclusion
 We propose that annotation of user preferences for text-to-image generation can be automated with analysis of user facial expression reaction to generated images, which can potentially increase the efficiency and scalability of human preference data collection.
-We present the FERGI dataset, comprising of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
-Specifically, AU4 activation is most consistently reflective of negative evaluation of the image, so we developed an AU4 valence score to represent the user’s evaluation of the image reflected by AU4 activation, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt.
-Firstly, the AU4 valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including the ensemble of them.
-Therefore, it can potentially be used to selectively annotate large-scale user preference data to be used for fine-tuning scoring models that do not rely on data of facial expression reaction.
-Secondly, the AU4 valence score can be integrated with pre-trained scoring models to improve their overall accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label data of large scale human preferences for fine-tuning text-to-image generative models.
-Thirdly, AU4 may be useful for understanding individual user preferences to create personalized image generation models.
-Additionally, we show that the AU4 valence score is the most reflective of the evaluation of the fidelity of the generated images among all models, which is complementary to the pre-trained scoring models, which better reflect the evaluation of image-text alignment.
-Finally, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction in this work.
+We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
+Specifically, we developed an AUcomb valence score to represent the user’s evaluation of the image reflected by AU4 and AU12 activations, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt.
+Firstly, the AUcomb valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including their ensemble.
+Therefore, it can potentially be used to selectively annotate large-scale user preference data used for fine-tuning scoring models.
+Note that automatic annotation will enable collection of vastly more human preference data from real users, so even if only a small portion are usable, its scale will still be much larger than manually annotated datasets.
+Secondly, the AUcomb valence score can be integrated with pre-trained scoring models to improve their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
+Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al.
+[75] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.44% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4.
+(see Tab.
+4 and Fig.
+6 of [75]).
+Thirdly, our AUcomb valence score may be useful for understanding individual user preferences to create personalized image generation models.
+In conclusion, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction.
 The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [53], image inpainting [21], and super-resolution [70].
@@
 Supplementary Material
-A Supplementary Visuals
-Figure S1: A flow chart summarizing the pipeline our paper proposes.
-The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated in the paper.
-Specifically, we developed a model and method for computing an AU4 valence score and integrating it with state-of-the-art text-to-image scoring models.
-We propose that the AU4 valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to improve/fine-tune the text-to-image generative models.
-Figure S2: An example of AU4 activation in response to a low-quality image generation.
-(a) The input text prompt.
-(b) The facial expression of the participant before seeing the generated image.
-(c) The generated image.
-(d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity).
-In the annotation survey, the participant gives an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she also ranked the image as the worst among the five images generated from the same prompt.
-To better illustrate the idea of our paper, we present Fig.
-S1 showing a flow chart summarizing the pipeline the paper proposed and Fig.
-S2 showing an example of a strong AU4 activation in response to a low-quality image generation.
-B Data Collection
+A Data Collection
 This section provides more details of the collection of the FERGI dataset not provided in Sec.
-2.
+3.
 Figs.
-S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
-B.1 Participants
+S1, S2, S3, S4, S5, S6 and S7 display screenshots of an example of one session in data collection.
+A.1 Participants
 39 participants [^1] The user IDs are up to 040 as shown in Tab.
@@
 Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community.
-We ended up with a dataset of 33 participants that is available upon request for research purposes.
-B.2 Prompt Creation
-There are two forms of sessions, structured input sessions and free-form input sessions.
-Most participants completed the same number of structured input sessions and free-input sessions.
+We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
+A.2 Prompt Creation
+There are two forms of sessions, structured input (SI) sessions and free-form input (FFI) sessions.
+Most participants completed the same number of structured input sessions and free-form input sessions.
 The number of sessions completed by each participant are shown in Tab.
@@
 internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
-User ID Number of Structured Input Sessions Number of Free-Form Input Sessions Number of Valid Images
+User ID Number of SI Sessions Number of FFI Sessions Number of Valid Images
 001 10 10 100
@@
 Total 288 288 2827
-Table S1: Information of the FERGI dataset.
-B.2.1 Structured Input
+Table S1: Information of the FERGI dataset. Numbers of structured input (SI) sessions, numbers of free-form input (FFI) sessions, and numbers of valid images for each participant are shown.
+A.2.1 Structured Input
 The purpose of having structured input sessions is to encourage creation of more diverse text prompts featuring different types of objects, interactions, settings, and styles.
@@
   • Animate Objects (Fig.
-2(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc.
+0(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc.
 Participants are required to enter the name and quantity of the object, with activity and a list of characteristics as optional inputs.
   • Inanimate Objects (Fig.
-2(b)): Any object that cannot do things, including plants, vehicles, furniture, etc.
+0(b)): Any object that cannot do things, including plants, vehicles, furniture, etc.
 Participants are required to enter the name and quantity of the object, with a list of characteristics as optional inputs.
   • Interactional Relations (Fig.
-2(c)): Relationship between an animated object and any other object (something an animate object does to another object).
+0(c)): Relationship between an animated object and any other object (something an animate object does to another object).
 An animate object can only be in a relation if its “activity” field is empty.
@@
   • Positional Relations (Fig.
-2(d)): Relationship between any two objects.
+0(d)): Relationship between any two objects.
 An animate object can only be in a relation if its “activity” field is empty.
@@
 Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs.
-S3 and S4.
-B.2.2 Free-Form Input
+S1 and S2.
+A.2.2 Free-Form Input
 The purpose of free-form input is to provide freedom and flexibility for participants to experiment with any prompts to complement the less flexible input style of structured input.
@@
 A screenshot of an example of the prompt creation page for a free-form input session is shown in Fig.
-S5.
-B.3 Image Presentation
-Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
-B.3.1 Image Annotation
+S3.
+A.3 Image Presentation
+Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S4 and S5.
+A.3.1 Image Annotation
 The annotation survey was adapted from the survey designed for ImageReward [75] and has the following components:
@@
 A screenshot of an example of the image annotation page (including the details of the questions) is shown in Fig. S6.
-B.3.2 Image Ranking
+A.3.2 Image Ranking
 After the participants finish annotating all 5 generated images, they will be asked to rank them from best to worst.
@@
 A screenshot of an example of the image ranking page is shown in Fig.
-S9.
+S7.
 (a) Animate objects
@@
 Figure S7: An example of an image ranking page.
-C AU Model Training
+B AU Model Training
 This section provides more details of the training of the AU model not provided in Sec. 4.
-C.1 Datasets
+B.1 Datasets
 The DISFA dataset [50] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total.
@@
 It contains facial video recordings of 9 participants’ posed and spontaneous facial expression with each frame being annotated with the same 12 AUs on a scale of 0 to 5.
-C.2 Video Preprocessing
+B.2 Video Preprocessing
 For each frame of the video, we first detect and crop the largest face in the image using MediaPipe [45] and then align the face with Pytorch Face Landmark Detection [7].
@@
 Similar preprocessing is also applied to the videos we analyze in the FERGI dataset.
-C.3 Model Training
+B.3 Model Training
 We use the neural network IR-50 [12] pre-trained on Glink360k [1] and fine-tune it on the DISFA and DISFA+ datasets [50, 48].
@@
 We train the model on all data from the DISFA and DISFA+ datasets [50, 48] for a total of 3 epochs using a single NVIDIA GeForce GTX 1080Ti 11G GPU.
-D Data Filtering
+C Data Filtering
 This section provides the details of frame exclusion not provided in Sec. 5.1.

### v3

@@
-FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction
+FERGI: Automatic Scoring of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction
 Abstract
@@
 However, the scalability of human feedback collection has been limited by its reliance on manual annotation.
-Therefore, we develop and test a method to automatically annotate user preferences from their spontaneous facial expression reaction to the generated images.
+Therefore, we develop and test a method to automatically score user preferences from their spontaneous facial expression reaction to the generated images.
 We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images.
-Specifically, AU4 (brow lowerer) is reflective of negative evaluations of the generated image whereas AU12 (lip corner puller) is reflective of positive evaluations.
-These can be useful in two ways.
-Firstly, we can automatically annotate user preferences between image pairs with substantial difference in these AU responses with an accuracy significantly outperforming state-of-the-art scoring models.
-Secondly, directly integrating the AU responses with the scoring models improves their consistency with human preferences.
-Finally, this method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks.
+We develop an FAU-Net (Facial Action Units Neural Network), which receives inputs from an AU estimation model, to automatically score user preferences for text-to-image generation based on their facial expression reactions, which is complementary to the pre-trained scoring models based on the input text prompts and generated images.
+Integrating our FAU-Net valence score with the pre-trained scoring models improves their consistency with human preferences.
+This method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks.
 The code is available at https://github.com/ShuangquanFeng/FERGI, and the dataset is also available at the same link for research purposes.
@@
 However, the scalability of human feedback collection has still been limited by its reliance on manual annotation.
-Therefore, we develop and test a system to automatically annotate user preferences for text-to-image generation from their spontaneous facial expression reaction to the generated images.
-This automatic annotation requires zero additional effort from the real users of text-to-image generators and is thus expected to increase the scalability of human preference feedback data collection.
+Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to automatically score human preferences of text-to-image generation based on the input text prompts and the generated images.
+In this study, we develop and test a system to automatically score user preferences from their spontaneous facial expression reaction to the generated images, which is complementary to the models based only on the texts and images.
+This automatic scoring requires zero additional effort from the real users of text-to-image generators.
 We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion (SD) v1.4 [62] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input.
-We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [20]), in the facial expression reaction videos with a pre-trained model and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images.
-We further propose a method of automatically annotating user preferences by detecting the activations of AU4 (brow lowerer, see Fig.
-5) and AU12 (lip corner puller, see Fig.
-6) in the facial expression reaction to the generated images.
-Figure 1: A flow chart summarizing our proposed pipeline.
-The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated.
-Specifically, we developed a model and method for computing an AUcomb valence score and integrating it with state-of-the-art text-to-image scoring models.
-We propose that the AUcomb valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to fine-tune the text-to-image generative models.
-To better illustrate the idea of our paper, we present Fig. 1 showing a flow chart summarizing the pipeline the paper proposed.
+We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [21]), in the facial expression reaction videos with a model we trained on external datasets and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images.
+We further propose a method of automatically scoring user preferences by detecting the activations of the AUs in the facial expression reaction to the generated images and feeding them into a neural network.
 2 Related Work
@@
 AU detection and estimation has attracted increasing interest [29, 65, 88, 87, 12, 45, 46, 73, 36, 37, 89, 49, 69, 70] and been used for facial emotion recognition [81, 83, 82, 80].
-Specifically related to our usage of AU4 and AU12 to infer user evaluations of generated images, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [72].
+For example, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [76].
 Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [26] and increases with negative stimuli, such as negative images [7, 6, 40, 39].
@@
 6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not permitting sharing of their video recordings for research.
-Therefore, we ended up with a dataset of 33 participants that is available upon request for research purposes at https://github.com/ShuangquanFeng/FERGI.
+This resulted in a dataset of 33 participants that will be made available for research.
 The collection and usage of human participant data in the dataset were approved by the Institutional Review Board (IRB) of UCSD.
@@
 A flow chart of the procedure of each session is shown in Fig.
-2.
+1.
 Details are explained below.
+Figure 1: Procedure of one session in data collection.
 3.2.1 Prompt Creation
 There are two types of sessions, structured input and free-form input.
-In both types, the participant can freely draft any text prompt as the input as long as it is not NSFW.
+In both types, the participant can freely draft any text prompt as long as it is not NSFW.
 The only difference between the two types lies in how the input text prompt is created.
@@
 (multi-choice question), and emotions felt when seeing the image (multi-choice question).
-After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server. That marks the end of the session.
+After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server ending the session.
 3.3 Data Preprocessing
@@
 The data associated with these images were excluded from further analysis, resulting in fewer than 5 valid images for some input text prompts.
-Figure 2: Procedure of one session in data collection.
 4 AU Model Training
@@
 In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop).
-Video frames are first preprocessed with face detection [45], face alignment [7], and a combination of histogram equalization and linear mapping [36] before being fed into the model for training, testing, or inference.
+(See Fig.
+S10 for a visual reference guide for the analyzed AUs.) Video frames are first preprocessed with face detection [48], face alignment [8], and a combination of histogram equalization and linear mapping [38] before being fed into the model for training, testing, or inference.
 For model training, we used the neural network IR-50 [13] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [53, 51].
 A single network learned the estimation of all labeled AUs with both regression and ordinal classification [55].
-More details of the training of the AU model is elaborated in the Supplementary Materials.
+More details of the training of the AU model is elaborated in the Appendix / Supplementary Materials.
+The comparison between our AU estimation model’s performance on the DISFA dataset and other state-of-the-art models, including OpenFace 2.0 [2], CCNN-IT [73], 2DC [47], SCC-Heatmap [23], and iARL [65] is shown in Tab.
+1.
+Our model’s Pearson’s $r$ (higher is better) is close to that of OpenFace 2.0; our model’s ICC(3,1) (higher is better) is the best among all models.
+Although our model’s MAE (lower is better) is less competitive among all models, we believe that ICC(3,1) is a better metric for evaluating the performance of AU estimation models on the DISFA dataset considering the high imbalance of the dataset.
+Table 1: Comparison of our AU estimation model’s performance on the DISFA dataset with other models
+AU 1 2 4 5 6 9 12 15 17 20 25 26 Avg.
+Pearson’s $r$ (higher is better)
+OpenFace 2.0 .64 .50 .70 .67 .59 .54 .85 .39 .49 .22 .85 .67 .59
+Our Model .53 .46 .74 .57 .56 .60 .85 .42 .41 .21 .94 .66 .58
+ICC(3,1) (higher is better)
+CCNN-IT .18 .15 .61 .07 .65 .55 .82 .44 .37 .28 .77 .54 .45
+2DC .70 .55 .69 .05 .59 .57 .88 .32 .10 .08 .90 .50 .50
+SCC-Heatmap .73 .44 .74 .06 .27 .51 .71 .04 .37 .04 .94 .78 .47
+iARL .13 .36 .68 .22 .56 .36 .86 .52 .37 .12 .96 .60 .48
+Our Model .53 .46 .74 .56 .55 .57 .85 .41 .40 .20 .93 .66 .57
+MAE (lower is better)
+CCNN-IT 87 .63 .86 .26 .73 .57 .55 .38 .57 .45 .81 .64 .61
+SCC-Heatmap .16 .16 .27 .03 .25 .13 .32 .15 .20 .09 .30 .32 .20
+iARL .30 .31 .52 .04 .36 .30 .31 .05 .33 .08 .29 .26 .26
+Our Model .37 .39 .50 .14 .37 .21 .33 .24 .47 .23 .32 .42 .33
 5 Facial Feature Extraction
@@
 Similar procedures were used to extract features from baseline clips and reaction clips.
-Further details are in the Supplemental Materials.
+Further details are in the Appendix / Supplementary Materials.
 5.1 Data Filtering
@@
 Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and thus for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [48]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [8]).
-Then, if a clip has more than $20\%$ of its total frames being excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
+If a clip has more than $20\%$ of its total frames excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
 5.2 AU Activation Value
-We compute an AU activation value $\alpha_{i}$ for each trained AUi (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
-  • Firstly, we estimate the intensity of each AU for each frame of the video clip with the AU model: $\hat{y}_{i,\mathrm{reg}}^{(1)},\hat{y}_{i,\mathrm{reg}}^{(2)},\dots$, where the superscript indicates the index of the frame.
-  • Secondly, we compute the moving window mean of the estimated AU intensities for every 0.1 seconds (3 frames for 30 FPS):
+We compute an AU activation value $\alpha_{i}$ for each trained AUi in each video clip. The AU activation value is computed as follows:
+  • Firstly, the intensity of each AU for each frame of the video clip is estimated with the AU model: $\hat{y}_{i,\mathrm{reg}}^{(1)},\hat{y}_{i,\mathrm{reg}}^{(2)},\dots$, where the superscript indicates the index of the frame.
+  • Secondly, a moving window mean of the estimated AU intensities for every 0.1 seconds (3 frames for 30 FPS) is computed:
     $\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)}=\frac{1}{3}\Sigma_{k^{\prime}=k}^{k+2}\hat{y}_{i,\mathrm{reg}}^{(k^{\prime})}$ (1)
     for all $k$ with defined $\hat{y}_{i,\mathrm{reg}}^{(k+1)}$ and $\hat{y}_{i,\mathrm{reg}}^{(k+2)}$ (as they may be undefined for frames at the end of the video clip or frames adjacent to excluded frames).
-  • Finally, we compute the AU activation of the clip as the difference between the maximum AU intensity within the video clip and the AU intensity of the first 0.1 seconds:
+  • Finally, the AU activation of the clip as the difference between the maximum AU intensity within the video clip and the AU intensity of the first 0.1 seconds is computed:
     $\alpha_{i}=\max_{k}(\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)})-\bar{\hat{y}}_{i,\mathrm{reg}}^{(1)}.$ (2)
@@
 […]
-(a) disappointed
-(b) satisfied
-(c) surprised
-(d) disgusted
-(e) amused
-(f) scared
+(a) surprised
+(b) disgusted
+(c) amused
+(d) scared
+(e) disappointed
+(f) satisfied
 Figure 3: Reported emotions highly correlated with AU activation values.
-The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?” Each subfigure shows the results for a different emotion (indicated in the subcaptions) and contains multiple pairs of boxplots representing the results for different AUs.
-The x-axis represents the indices of the AUs while the y-axis represents the AU activation values in the reaction clip of each image.
-The parenthesized numbers in the legends represent the number of responses of that type.
+The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?”.
+Each subfigure shows results for a different emotion (indicated in the captions) and contains multiple pairs of boxplots representing the results for different AUs.
+The indices of the AUs are along the x-axis, while the y-axis represents the AU activation values in the reaction clip of each image.
+The parenthesized numbers in the legends give the number of responses of that type.
 In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion.
-In each boxplot, the bottom of the box represents the first quartile (25th percentile) of the AU activation values, the top of the box represents the third quartile (75th percentile), and the line in the middle of the box represents the median.
+In each boxplot, the bottom/top of the box represents the first/third quartile (25th/75th percentile) of the AU activation values, and the line in the middle of the box represents the median.
 We started with analyzing statistical relationships between the AU activation values and answers in the annotation survey for each image. The primary results are as follows:
@@
   • Emotions.
-For the question “Did you feel any of the following emotions when you saw the image?”, the participant can independently select 0 to 6 answers from the six options: disappointed, satisfied, surprised, disgusted, amused, and scared.
+For the question “Did you feel any of the following emotions when you saw the image?”, the participant can independently select 0 to 6 answers from the six options: surprised, disgusted, amused, scared, disappointed, and satisfied.
 Each of the reported emotions was found to be significantly correlated with the activation values of multiple AUs.
+    • Surprised.
+“Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig.
+3(a) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
+    • Disgusted.
+“Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig.
+3(b) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
+    • Amused.
+“Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig.
+3(c) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
+    • Scared.
+“Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig.
+3(d) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
     • Disappointed.
 “Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig.
-3(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
+3(e) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
     • Satisfied.
 “Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig.
-3(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
-    • Surprised.
-“Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig.
-3(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
-    • Disgusted.
-“Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig.
-3(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
-    • Amused.
-“Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig.
-3(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
-    • Scared.
-“Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig.
-3(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
-Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$, with 12 representing the 12 investigated AUs, 2 representing the two tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 representing the 6 tests for 6 emotions.
-6.2 Automatic Annotation of User Preferences between Image Pairs Using AU4 and AU12
-Figure 5: An example of AU4 activation in response to a low-quality image generation.
+3(f) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
+Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$ (12 investigated AUs, 2 tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 emotions).
+6.2 Automatic Annotation of User Preferences between Image Pairs
+6.2.1 Pre-trained Scoring Models
+Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to estimate human preferences of text-to-image generation based on the input text prompts and the generated images.
+Among these baseline scoring models, ImageReward, PickScore, and HPS v2 were all trained specifically on datasets of human preferences of text-to-image generation and are supposed to outperform the other models [79, 35, 77].
+To introduce a potentially more competitive baseline score, we further combine them to form an ensemble baseline score
+$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$ (3)
+subject $\text{subject to}\quad w_{\mathrm{IR}}+w_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}$ $=1,$ (4)
+where $s_{\mathrm{ens}}$ represents the ensemble baseline score, $s_{\mathrm{IR}}$, $s_{\mathrm{Pick}}$, and $s_{\mathrm{HPSv2}}$ represent the ImageReward score, PickScore, and HPS v2 score respectively, and $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$ represent their respective weights.
+A leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction.
+Since the scales of the scores are different, they were each z-scored using their mean and standard deviation on the training set.
+We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
+6.2.2 FAU-Net Valence Score
+We aim to use the estimated AU activation values to automatically score user preferences between images generated from the same input text prompt and propose that this scoring more directly reflects the user’s reaction to the generated image and is complementary to the pre-trained scoring models.
+To achieve this goal, we train an FAU-Net (Facial Action Units Neural Network) to do automatic scoring.
+In the FAU-Net, the estimated activation values of the 12 investigated AUs are used as the inputs; each activation value is first preprocessed with a linear transformation with trained parameters and sigmoid function ($\sigma(a\cdot\alpha+b)$ with the initializations of $a=1$ and $b=-1$ for encouraging transformations of positive correlation for better interpretability); then the preprocessed AU activation values are used as inputs to a hidden layer with 16 neurons which feeds into a single ouput neuron representing the AU evaluation score of the image; this output unit receives inputs from both the 16 neurons of the hidden layer and the 12 preprocessed AU activation values.
+The sigmoid function is used as the activation function for all neurons.
+The network is trained with a RankNet loss function using the reaction clips and ranking order given by the participant for image pairs generated from the same text prompts and an L1 regularization with a coefficient of $10^{-4}$.
+For model training, we use the Adam optimizer with an initial learning rate of $10^{-3}$ and a single batch for all data to train the neural network for 300 epochs using a single NVIDIA RTX A6000 48GB GDDR6 GPU for about 15 minutes.
+Our FAU-Net is trained with an LOPO procedure and outputs an FAU-Net valence score $s_{\mathrm{FAU}}$ reflecting the evaluation of the generated image derived from the facial expression reaction.
+Note that the AU estimation model is frozen and not trained with the FERGI data.
+6.2.3 Integrate FAU-Net Valence Score with Pre-trained Scoring Models
+We integrate our FAU-Net valence score and the pre-trained scoring model with a simple linear combination:
+$s_{m}+a_{m}s_{\mathrm{FAU}},$ (5)
+where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score from model $m$, and $a_{m}$ the weight for integrating $m$ and the FAU-Net valence score.
+We separately investigated the effect of integrating the FAU-Net valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble.
+Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); the FAU-Net weights and the weightings $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, $w_{\mathrm{HPSv2}}$ computed in Eq.
+3 are not modified at this stage.
+Table 2: Accuracy of image preference prediction for a total of 3734 image pairs.
+The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the FAU-Net valence score.
+Model Ind. Acc. Acc. w/ AUs
+CLIP Score 57.74 /
+Aesthetic Score 50.32 /
+BLIP Score 50.05 /
+FAU-Net Valence Score 59.35 /
+ImageReward Score 60.47 65.21
+PickScore 63.28 66.28
+HPS V2 Score 62.00 66.87
+Ensemble Baseline Score 65.80 68.64
+(a) Selected by
+ImageReward score
+(b) Selected by
+PickScore
+(c) Selected by
+HPS V2 Score
+(d) Selected by
+FAU-Net Valence Score
+Figure 4: FAU-Net valence score is complementary to the other pre-trained scoring models.
+Each subfigure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (given in the subcaption label), with the x-axis representing the proportion of data selected (image pairs are selected based on the largest absolute difference of scores for the two images) and the y-axis representing the annotation accuracy within the selected subset.
+Each model is good at estimating the image pairs that it will perform well on as evidenced by increased accuracy for a smaller selected proportion.
+Note the blue, green, and orange curves are almost on top of each other.
+(a) ${\color[rgb]{1,0,0}w_{h(1)-o}\approx-0.58}$
+(b) ${\color[rgb]{1,0,0}w_{h(2)-o}\approx-0.53}$
+(c) ${\color[rgb]{1,0,0}w_{h(3)-o}\approx-0.50}$
+(d) ${\color[rgb]{1,0,0}w_{h(4)-o}\approx-0.49}$
+(e) ${\color[rgb]{1,0,0}w_{h(5)-o}\approx-0.46}$
+(f) ${\color[rgb]{1,0,0}w_{h(6)-o}\approx-0.45}$
+(g) ${\color[rgb]{1,0,0}w_{h(7)-o}\approx-0.45}$
+(h) ${\color[rgb]{1,0,0}w_{h(8)-o}\approx-0.42}$
+(i) ${\color[rgb]{1,0,0}w_{h(9)-o}\approx-0.37}$
+(j) ${\color[rgb]{1,0,0}w_{h(10)-o}\approx-0.22}$
+(k) ${\color[rgb]{1,0,0}w_{h(11)-o}\approx-0.13}$
+(l) ${\color[rgb]{1,0,0}w_{h(12)-o}\approx-0.12}$
+(m) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
+(n) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
+(o) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
+(p) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
+Figure 5: Weights of hidden nodes.
+Each subfigure shows the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption illustrating the weight from the hidden node to the output node (red/green for negative/positive weights).
+The hidden nodes are ordered by the weights from them to the output node in this figure.
+Figure 6: Weights of the output node. The weights from the input preprocessed activation values of the 12 AUs to the output node in the FAU-Net.
+The results are shown in Tab.
+2.
+Integration with FAU-Net valence score improves the performance of all four baseline models.
+Specifically, by further integrating all three pre-trained human preference scoring models together with the FAU-Net valence score, we achieve the highest accuracy of 68.64%, outperforming the ensemble of the three models by a margin of 2.84% ($p<0.01$ with a two-proportion z-test).
+6.2.4 Complementarity between FAU-Net and Pre-trained Scoring Models
+The complementarity between FAU-Net and pre-trained scoring models is demonstrated in Fig.
+4.
+The figure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (image pairs are selected based on the largest absolute difference scores for the two images given by the associated scoring model).
+We can clearly see that the three pre-trained scoring models are highly correlated showing similar performance curves (almost overlapping).
+In contrast, our FAU-Net valence score selects different image pairs, explaining why the FAU-Net valence scores combine synergistically with the other scores.
+6.2.5 FAU-Net Analysis
+We analyze the weights of the trained FAU-Net to better understand the model.
+Figure 5 shows the weights from the input preprocessed AU activation values to the nodes in the hidden layer (given that the preprocessings of activation values are still all increasing functions after model training) with the subcaptions illustrating the weight from the corresponding hidden node to the output node.
+Generally, for hidden nodes positively affecting the output score, they receive positive weights from the activations of AU2, AU6, AU12, AU17, AU20, and AU26 and negative weights from the activations of AU4, AU5, AU9, AU17, and AU25, and for hidden nodes negatively affecting the output score, they receive weights of the opposite signs from the AUs.
+Similar trends are observed for the weights from the input preprocessed AU activation values and the output node as shown in Fig.
+6.
+7 Discussion and Conclusion
+We propose that user facial expression reactions to text-to-image generation can be used to score user preferences for the generated images, which is complementary to existing scoring models which estimate human preferences based only on the input text prompts and generated images.
+We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
+Specifically, we developed an FAU-Net that receives the detected AU activations in the facial expression reaction as inputs and outputs an FAU-Net valence score as the estimation of the user preference.
+Integrating the FAU-Net valence score with pre-trained scoring models improves their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
+Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al.
+[79] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.47% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4.
+(see Tab.
+4 and Fig.
+6 of [79]).
+We also demonstrated the complementarity between our FAU-Net and the pre-trained scoring models, explaining why integrating them effectively improves the annotation accuracy.
+As they are respectively better at annotating different sets of image pairs, it is also potentially feasible to use them to fine-tune each other on data where one model gives larger absolute difference of scores for the two images.
+In conclusion, we have demonstrated the feasibility of automatically scoring user preferences for image generation from facial expression reaction.
+The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [57], image inpainting [22], and super-resolution [74].
+8 Acknowledgments
+We thank Xiaojing Xu and Yuan Tang for helpful prior work and Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions.
+We are grateful for support from NSF IIS 1817226 and seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion as well as hardware funding from NVIDIA, Adobe, and Sony.
+Supplementary Material
+A Supplementary Visuals
+To better illustrate the idea of our paper, we present Fig.
+S1 showing an example of a strong AU4 activation in response to a low-quality image generation and Fig.
+S2 an example of a strong AU12 activation in response to a high-quality image generation.
+Figure S1: An example of AU4 activation in response to a low-quality image generation.
 (a) The input text prompt.
@@
 In the annotation survey, the participant gave an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she ranked the image as the worst among the five images generated from the same prompt.
-Figure 6: An example of AU12 activation in response to a high-quality image generation.
+Figure S2: An example of AU12 activation in response to a high-quality image generation.
 (a) The input text prompt.
@@
 In the annotation survey, the participant gave ratings of 7 for each of image-text alignment, fidelity, and overall rating; he did not identify any issues with the image; he reported feeling “satisfied”, “surprised”, and “amused” when seeing the image; he ranked the image as the best among the five images generated from the same prompt.
-We aim to use the estimated AU activation values to automatically annotate user preferences between images generated from the same input text prompt.
-We have observed that the activations of AU2, AU4, AU9, and AU12 are significantly correlated with the the raw overall ratings in the previous subsection.
-After comparing combinations between them, we developed the following pipeline of using AU4 and AU12 to do automatic annotation (examples of AU4 and AU12 responses are shown in Figs.
-5 and 6).
-(Inclusion of AU2 and AU9 did not help.) We also did not find a way to effectively incorporate the information conveyed by AUs that correlate only with the extremity of the ratings.
-Collection of a much larger-scale dataset in the future might enable effective incorporation of such information with a more complex model.
-6.2.1 Participant Exclusion
-While estimations of AU12 are generally reliable and stable, multiple factors may result in unreliable AU4 estimation from a small number of people by our AU model.
-We excluded 3 participants with unreliable AU4 estimation due to bangs covering the area between the eyebrows (2 participants) or highly unstable camera angles (1 participant).
-6.2.2 AUcomb Valence Score
-The AU4 activation value $\alpha_{4}$ and AU12 activation value $\alpha_{12}$ are good indicators of the user’s evaluation of the generated image, but the raw values might not linearly reflect the associated evaluation.
-Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$ and an AU12 valence score $s_{\mathrm{AU12}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image.
-We define the $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ and $s_{\mathrm{AU12}}$ as the positive exponential decay function of AU12 activation value $\alpha_{12}$ as follows:
-$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}})$ (3)
-and
-$s_{\mathrm{AU12}}=(1-e^{-k\cdot\alpha_{12}}),$ (4)
-where $k$ is an exponential decay coefficient to fit. We further define the final AUcomb valence score $s_{\mathrm{AUcomb}}$ as the linear combination of them:
-$s_{\mathrm{AUcomb}}=w_{\mathrm{AU4}}s_{\mathrm{AU4}}+w_{\mathrm{AU12}}s_{\mathrm{AU12}}$ (5)
-subject $\text{subject to}\quad w_{\mathrm{AU4}}+w_{\mathrm{AU12}}$ $=1,$ (6)
-where $w_{\mathrm{AU4}}$ and $w_{\mathrm{AU12}}$ represent the weights for AU4 and AU12 respectively.
-6.2.3 Baseline Scores
-To compare the performance of our AU valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [41], ImageReward score [75], PickScore [33], and HPS v2 score [73] as baselines.
-Among these baseline scoring models, ImageReward, PickScore, and HPS v2 were all trained specifically on datasets of human preferences of text-to-image generation and are supposed to outperform the other models [75, 33, 73].
-To introduce a potentially more competitive baseline score, we further combine them to form an ensemble baseline score
-$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$ (7)
-subject $\text{subject to}\quad w_{\mathrm{IR}}+w_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}$ $=1,$ (8)
-where $s_{\mathrm{ens}}$ represents the ensemble baseline score, $s_{\mathrm{IR}}$, $s_{\mathrm{Pick}}$, and $s_{\mathrm{HPSv2}}$ represent the ImageReward score, PickScore, and HPS v2 score respectively, and $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$ represent their respective weights.
-To be consistent with the evaluation of AUcomb valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction.
-Since the scales of the scores are different, they were each z-scored using their mean and standard deviation on the training set.
-We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
-6.2.4 AUcomb Valence Score: Independent Preference Prediction for Selected Image Pairs
-We first evaluate the AUcomb valence score on its ability to independently predict user preferences between pairs of images generated by the same text prompt.
-Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data on a total of 3534 image pairs from the 30 included participants.
-Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AUcomb valence scores may not be informative for inferring preference between the two associated images.
-Therefore, we apply a threshold $d$ for making predictions: only when the difference between the AUcomb valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AUcomb valence score; otherwise, we abstain from making a prediction.
-There are 3 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eqs.
-3 and 4), the weight for AU4 $w_{\mathrm{AU4}}$ (as in Eq.
-5; $w_{\mathrm{AU12}}$ is then automatically set by Eq.
-6) and the difference threshold $d$.
-We fit these three parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$, from 0.0 to 1.0 with a granularity of 0.1 for $w_{\mathrm{AU4}}$, and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
-$n_{\mathrm{correct}}-2n_{\mathrm{incorrect}},$ (9)
-where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the number of correct predictions and incorrect predictions on image pair preference.
-Table 1: Accuracy of image preference prediction on the 19.38% of all data (685 out of 3534 pairs) as selected by the AUcomb valence model.
-(Image pairs with AUcomb valence score difference below the trained difference threshold are not included.)
-Model Acc.
-CLIP Score 58.25
-Aesthetic Score 51.39
-BLIP Score 49.78
-ImageReward Score 62.19
-PickScore 61.31
-HPS V2 Score 62.04
-Ensemble Baseline Score 66.57
-AUcomb Valence Score 72.70
-The results are shown in Tab.
-1.
-Based on the trained difference thresholds, the AUcomb valence model makes predictions on 19.38% of the image preference data (685 out of 3534 pairs).
-For the selected image pairs, our AUcomb valence score has the best accuracy of 72.70%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 6.13%.
-Our results have high stability.
-The parameter values fit in the LOPO procedure were generally consistent.
-We obtained $k=0.5$, $w_{\mathrm{AU4}}=0.5$, $w_{\mathrm{AU12}}=0.5$, and $d=0.26$ for 20 out of the 30 total training sets.
-6.2.5 Integrate AUcomb Valence Score with Pre-trained Scoring Models
-AUcomb valence score has a great performance of 72.70% on the subset of the image pairs with distinguishable AUcomb valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 58.23%, still significantly better than chance (two-sided binomial test: $p<5\times 10^{-22}$) but lower than that of the ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab.
-2).
-However, besides making independent predictions, the AUcomb valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data.
-We integrate the pre-trained scoring models with AUcomb valence score with a simple linear combination:
-$s_{m}+a_{m}s_{\mathrm{AUcomb}},$ (10)
-where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score from model $m$, and $a_{m}$ the weight for integrating $m$ and the AUcomb valence score.
-We separately investigated the effect of integrating the AUcomb valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble.
-Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); for the weights of the ensemble baseline score (Eq.
-7), exponential decay coefficient for the AU4 and AU12 valence scores (Eqs.
-3 and 4), and the weighting between AU4 and AU12, we directly used the values fit previously.
-Table 2: Accuracy of image preference prediction on all data.
-The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AUcomb valence score.
-Model Ind. Acc. Acc. w/ AUs
-CLIP Score 57.50 /
-Aesthetic Score 49.92 /
-BLIP Score 49.92 /
-AUcomb Valence Score 58.23 /
-ImageReward Score 60.44 63.67
-PickScore 63.36 65.42
-HPS V2 Score 61.94 65.53
-Ensemble Baseline Score 65.79 67.49
-The results are shown in Tab.
-2.
-Integration with AUcomb valence score improves the performance of all of the four baseline models.
-Specifically, by further integrating all three pre-trained human preference scoring models together with the AUcomb valence score, we achieve the highest accuracy of 67.49%, outperforming the ensemble of the three models by a margin of 1.70%.
-7 Discussion and Conclusion
-We propose that annotation of user preferences for text-to-image generation can be automated with analysis of user facial expression reaction to generated images, which can potentially increase the efficiency and scalability of human preference data collection.
-We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
-Specifically, we developed an AUcomb valence score to represent the user’s evaluation of the image reflected by AU4 and AU12 activations, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt.
-Firstly, the AUcomb valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including their ensemble.
-Therefore, it can potentially be used to selectively annotate large-scale user preference data used for fine-tuning scoring models.
-Note that automatic annotation will enable collection of vastly more human preference data from real users, so even if only a small portion are usable, its scale will still be much larger than manually annotated datasets.
-Secondly, the AUcomb valence score can be integrated with pre-trained scoring models to improve their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
-Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al.
-[75] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.44% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4.
-(see Tab.
-4 and Fig.
-6 of [75]).
-Thirdly, our AUcomb valence score may be useful for understanding individual user preferences to create personalized image generation models.
-In conclusion, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction.
-The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [53], image inpainting [21], and super-resolution [70].
-8 Acknowledgments
-We thank Xiaojing Xu and Yuan Tang for helpful prior work and Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions.
-We are grateful for support from NSF IIS 1817226 and seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion as well as hardware funding from NVIDIA, Adobe, and Sony.
-Supplementary Material
-A Data Collection
+B Data Collection
 This section provides more details of the collection of the FERGI dataset not provided in Sec.
@@
 Figs.
-S1, S2, S3, S4, S5, S6 and S7 display screenshots of an example of one session in data collection.
-A.1 Participants
+S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
+B.1 Participants
 39 participants [^1] The user IDs are up to 040 as shown in Tab.
@@
 were recruited from the SONA system of University of California San Diego (UCSD).
+and received academic credits as compensation.
 They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers.
@@
 We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
-A.2 Prompt Creation
+B.2 Prompt Creation
 There are two forms of sessions, structured input (SI) sessions and free-form input (FFI) sessions.
@@
 Table S1: Information of the FERGI dataset. Numbers of structured input (SI) sessions, numbers of free-form input (FFI) sessions, and numbers of valid images for each participant are shown.
-A.2.1 Structured Input
+B.2.1 Structured Input
 The purpose of having structured input sessions is to encourage creation of more diverse text prompts featuring different types of objects, interactions, settings, and styles.
@@
   • Animate Objects (Fig.
-0(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc.
+3(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc.
 Participants are required to enter the name and quantity of the object, with activity and a list of characteristics as optional inputs.
   • Inanimate Objects (Fig.
-0(b)): Any object that cannot do things, including plants, vehicles, furniture, etc.
+3(b)): Any object that cannot do things, including plants, vehicles, furniture, etc.
 Participants are required to enter the name and quantity of the object, with a list of characteristics as optional inputs.
   • Interactional Relations (Fig.
-0(c)): Relationship between an animated object and any other object (something an animate object does to another object).
+3(c)): Relationship between an animated object and any other object (something an animate object does to another object).
 An animate object can only be in a relation if its “activity” field is empty.
@@
   • Positional Relations (Fig.
-0(d)): Relationship between any two objects.
+3(d)): Relationship between any two objects.
 An animate object can only be in a relation if its “activity” field is empty.
@@
 Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs.
-S1 and S2.
-A.2.2 Free-Form Input
+S3 and S4.
+B.2.2 Free-Form Input
 The purpose of free-form input is to provide freedom and flexibility for participants to experiment with any prompts to complement the less flexible input style of structured input.
@@
 A screenshot of an example of the prompt creation page for a free-form input session is shown in Fig.
-S3.
-A.3 Image Presentation
-Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S4 and S5.
-A.3.1 Image Annotation
+S5.
+B.3 Image Presentation
+Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
+B.3.1 Image Annotation
 The annotation survey was adapted from the survey designed for ImageReward [79] and has the following components:
@@
 A screenshot of an example of the image annotation page (including the details of the questions) is shown in Fig. S8.
-A.3.2 Image Ranking
+B.3.2 Image Ranking
 After the participants finish annotating all 5 generated images, they will be asked to rank them from best to worst.
@@
 A screenshot of an example of the image ranking page is shown in Fig.
-S7.
+S9.
 (a) Animate objects
@@
 Figure S9: An example of an image ranking page.
-B AU Model Training
+C AU Model Training
 This section provides more details of the training of the AU model not provided in Sec. 4.
-B.1 Datasets
+C.1 Datasets
 The DISFA dataset [53] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total.
-Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5222See https://melindaozel.com/facs-cheat-sheet/ for a visual reference guide for FACS.
+Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5.
+See Fig.
+S10 for a visual reference guide for the analyzed AUs.
+Figure S10: A visual reference guide for the analyzed AUs extracted from [56].
 The DISFA+ dataset [51] is an extension of the DISFA dataset [53].
 It contains facial video recordings of 9 participants’ posed and spontaneous facial expression with each frame being annotated with the same 12 AUs on a scale of 0 to 5.
-B.2 Video Preprocessing
+C.2 Video Preprocessing
 For each frame of the video, we first detect and crop the largest face in the image using MediaPipe [48] and then align the face with Pytorch Face Landmark Detection [8].
@@
 Similar preprocessing is also applied to the videos we analyze in the FERGI dataset.
-B.3 Model Training
+C.3 Model Training
 We use the neural network IR-50 [13] pre-trained on Glink360k [1] and fine-tune it on the DISFA and DISFA+ datasets [53, 51].
@@
 For model training, we employ the Adam optimizer with an initial learning rate of $10^{-4}$ for parameters of the last layer and an $10^{-5}$ for other parameters, a weight decay of $5\times 10^{-4}$, and a batch size of 64.
-We train the model on all data from the DISFA and DISFA+ datasets [50, 48] for a total of 3 epochs using a single NVIDIA GeForce GTX 1080Ti 11G GPU.
-C Data Filtering
+We train the model on all data from the DISFA and DISFA+ datasets [53, 51] for a total of 3 epochs using a single NVIDIA GeForce GTX 1080Ti 11G GPU for about 1 hour.
+D Data Filtering
 This section provides the details of frame exclusion not provided in Sec. 5.1.

### v4

@@
 In this study, we develop and test a system to automatically score user preferences from their spontaneous facial expression reaction to the generated images, which is complementary to the models based only on the texts and images.
-This automatic scoring requires zero additional effort from the real users of text-to-image generators.
+This automatic scoring requires zero additional effort from the real users of text-to-image generative models.
 We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion (SD) v1.4 [63] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input.
@@
 More recent advancements in DMs [54, 63, 62, 59, 64, 14] have achieved great success in generating high-quality images and attracted widespread public attention.
-As traditional evaluation metrics like Inception Score (IS) [3], Fréchet Inception Distance (FID) [27], and CLIP score [59] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [78], ImageReward [79], PickScore [35], and Human Preference Score v2 (HPS v2) [77], and demonstrated the effectiveness of using them to improve the text-to-image generative models [79, 35, 77, 24].
+As traditional evaluation metrics like Inception Score (IS) [2], Fréchet Inception Distance (FID) [27], and CLIP score [60] failed to comprehensively capture human preferences for text-to-image generation, researchers have proposed to specifically train human preference scoring models based on large-scale human feedback datasets, such as Human Preference Score (HPS) [79], ImageReward [80], PickScore [35], and Human Preference Score v2 (HPS v2) [78], and demonstrated their effectiveness to improve text-to-image generative models [80, 35, 78, 24].
 Human Feedback Datasets.
@@
 As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [20, 19, 50], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [21].
-AU detection and estimation has attracted increasing interest [29, 65, 88, 87, 12, 45, 46, 73, 36, 37, 89, 49, 69, 70] and been used for facial emotion recognition [81, 83, 82, 80].
+AU detection and estimation has attracted increasing interest [29, 66, 91, 90, 12, 45, 46, 74, 36, 37, 92, 49, 70, 71] and been used for facial emotion recognition [83, 85, 84, 82, 81].
 For example, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [77].
@@
 39 participants were recruited from the SONA system of University of California San Diego (UCSD) and completed the study asynchronously on personal computers.
-6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not permitting sharing of their video recordings for research.
-This resulted in a dataset of 33 participants that will be made available for research.
-The collection and usage of human participant data in the dataset were approved by the Institutional Review Board (IRB) of UCSD.
+6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not allowing sharing of their video for research.
+Data collection and use were approved by the Institutional Review Board (IRB) of UCSD.
 3.2 Data Collection Procedure
@@
 In each session, the participant drafted one text prompt and viewed 5 images generated from the input text prompt.
-A flow chart of the procedure of each session is shown in Fig.
-1.
+A flow chart of the procedure of each session is shown in Figure 1.
 Details are explained below.
 Figure 1: Procedure of one session in data collection.
+Firstly, the participant creates an input text prompt.
+Secondly, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately; at the same time, 5 images are generated from the text prompt using Stable Diffusion v1.4.
+Then, for each image, the participant goes through a 5-second baseline phase, a 5-second image presentation phase, and an image annotation phase with no time restriction.
+(During the baseline phase, the participant’s baseline facial expression is recorded, and during the image presentation phase, the participant’s facial expression reaction to the generated image is recorded.) Finally, after all 5 images have been presented and annotated, the participant ranks the 5 generated images from best to worst.
+A copy of this figure with additional references to the screenshots of each stage is in the supplemental material.
 3.2.1 Prompt Creation
@@
 In both types, the participant can freely draft any text prompt as long as it is not NSFW.
-The only difference between the two types lies in how the input text prompt is created.
-In a structured input session, the participant fills out a form to separately specify different elements desired in the images to generate: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords.
-Based on the contents of the form, a text prompt is then automatically generated by the Large Language Model (LLM) OpenAI gpt3.5-turbo.
-The participant can freely adjust the generated prompt as the final input text prompt.
+The only difference lies in how the input text prompt is created.
+In a structured input session, the participant fills out a form to separately specify different elements desired in the generated images: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords.
+Based on the contents of the form, an initial text prompt is automatically generated by the Large Language Model (LLM) OpenAI GPT-3.5 Turbo which the participant can freely adjust to obtain the final input text prompt.
 In a free-form input, the participant directly enters the entire final input text prompt.
@@
 After the input text prompt is finalized, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately.
-The webcam is then used to record throughout the rest of the session.
+The webcam is used to record throughout the rest of the session.
 The configuration of the raw videos varied based on the participant’s device and browser.
-Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos were standardized to 30 FPS afterwards.
+However, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos were standardized to 30 FPS afterwards.
 3.2.3 Image Presentation, Annotation, and Ranking
@@
 The image presentation and annotation followed the webcam preview.
-The detailed procedures for the presentation and annotation of each generated image are as follows:
+The details are as follows:
   • Baseline (5 seconds): The website displays the “preparing image” status text at the center of the page along with the input text prompt.
-No specific facial-expression-eliciting event is expected to happen during this period.
+No specific facial-expression-eliciting event is expected during this period.
   • Present (5 seconds): The generated image is presented at the center of the webpage along with the input text prompt.
@@
 The participant annotates the image manually by filling out the survey during this interval.
-The survey was adapted from the survey designed for ImageReward [79] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (later fixed) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis.
-(multi-choice question), and emotions felt when seeing the image (multi-choice question).
+The survey was adapted from the survey designed for ImageReward [80] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (later fixed) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so answers for this question should not be used in analysis.
+(multi-answer question), and emotions felt when seeing the image (multi-answer question).
 After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server ending the session.
@@
 4 AU Model Training
-We analyze the facial expression reaction of the users by estimating the activation of their AUs.
+We analyze the facial expression reaction of the users by estimating their AU activations.
 We trained our AU estimation model on the DISFA [53, 52] and DISFA+ [51] datasets.
@@
 In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop).
-(See Fig.
-S10 for a visual reference guide for the analyzed AUs.) Video frames are first preprocessed with face detection [48], face alignment [8], and a combination of histogram equalization and linear mapping [38] before being fed into the model for training, testing, or inference.
+(See the supplemental material for a visual reference guide for the analyzed AUs.) Video frames are first preprocessed with face detection [48], face alignment [8], and a combination of histogram equalization and linear mapping [38] before being fed into the model for training, testing, or inference.
 For model training, we used the neural network IR-50 [13] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [53, 51].
 A single network learned the estimation of all labeled AUs with both regression and ordinal classification [56].
-More details of the training of the AU model is elaborated in the Appendix / Supplementary Materials.
-The comparison between our AU estimation model’s performance on the DISFA dataset and other state-of-the-art models, including OpenFace 2.0 [2], CCNN-IT [73], 2DC [47], SCC-Heatmap [23], and iARL [65] is shown in Tab.
-1.
-Our model’s Pearson’s $r$ (higher is better) is close to that of OpenFace 2.0; our model’s ICC(3,1) (higher is better) is the best among all models.
+More details of the training of the AU model are elaborated in the supplemental material.
+The comparison between our AU estimation model’s performance on the DISFA dataset and other state-of-the-art models is shown in Table 1.
+Our model’s ICC(3,1) (higher is better) is the best among all models.
 Although our model’s MAE (lower is better) is less competitive among all models, we believe that ICC(3,1) is a better metric for evaluating the performance of AU estimation models on the DISFA dataset considering the high imbalance of the dataset.
+Also, ICC(3,1) is specifically more important in our application here because we care about the change of facial expression in reaction to the generated images instead of its absolute intensity.
 Table 1: Comparison of our AU estimation model’s performance on the DISFA dataset with other models
-AU 1 2 4 5 6 9 12 15 17 20 25 26 Avg.
-Pearson’s $r$ (higher is better)
-OpenFace 2.0 .64 .50 .70 .67 .59 .54 .85 .39 .49 .22 .85 .67 .59
-Our Model .53 .46 .74 .57 .56 .60 .85 .42 .41 .21 .94 .66 .58
-ICC(3,1) (higher is better)
-CCNN-IT .18 .15 .61 .07 .65 .55 .82 .44 .37 .28 .77 .54 .45
-2DC .70 .55 .69 .05 .59 .57 .88 .32 .10 .08 .90 .50 .50
-SCC-Heatmap .73 .44 .74 .06 .27 .51 .71 .04 .37 .04 .94 .78 .47
-iARL .13 .36 .68 .22 .56 .36 .86 .52 .37 .12 .96 .60 .48
-Our Model .53 .46 .74 .56 .55 .57 .85 .41 .40 .20 .93 .66 .57
-MAE (lower is better)
-CCNN-IT 87 .63 .86 .26 .73 .57 .55 .38 .57 .45 .81 .64 .61
-SCC-Heatmap .16 .16 .27 .03 .25 .13 .32 .15 .20 .09 .30 .32 .20
-iARL .30 .31 .52 .04 .36 .30 .31 .05 .33 .08 .29 .26 .26
-Our Model .37 .39 .50 .14 .37 .21 .33 .24 .47 .23 .32 .42 .33
+Metric Method AU Average
+1 2 4 5 6 9 12 15 17 20 25 26
+ICC(3,1$\uparrow$ CCNN-IT [74] .18 .15 .61 .07 .65 .55 .82 .44 .37 .28 .77 .54 .45
+2DC [47] .70 .55 .69 .05 .59 .57 .88 .32 .10 .08 .90 .50 .50
+SCC-Heatmap [23] .73 .44 .74 .06 .27 .51 .71 .04 .37 .04 .94 .78 .47
+iARL [66] .13 .36 .68 .22 .56 .36 .86 .52 .37 .12 .96 .60 .48
+[…]
+[…]
+SCC-Heatmap [23] .16 .16 .27 .03 .25 .13 .32 .15 .20 .09 .30 .32 .20
+iARL [66] .30 .31 .52 .04 .36 .30 .31 .05 .33 .08 .29 .26 .26
+[…]
 5 Facial Feature Extraction
@@
 Similar procedures were used to extract features from baseline clips and reaction clips.
-Further details are in the Appendix / Supplementary Materials.
+Further details are in the supplemental material.
 5.1 Data Filtering
@@
   • Secondly, a moving window mean of the estimated AU intensities for every 0.1 seconds (3 frames for 30 FPS) is computed:
-    $\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)}=\frac{1}{3}\Sigma_{k^{\prime}=k}^{k+2}\hat{y}_{i,\mathrm{reg}}^{(k^{\prime})}$ (1)
+    $\bar{\hat{y}}_{i,\mathrm{reg}}^{(k)}=\frac{1}{3}\sum_{k^{\prime}=k}^{k+2}\hat{y}_{i,\mathrm{reg}}^{(k^{\prime})}$ (1)
     for all $k$ with defined $\hat{y}_{i,\mathrm{reg}}^{(k+1)}$ and $\hat{y}_{i,\mathrm{reg}}^{(k+2)}$ (as they may be undefined for frames at the end of the video clip or frames adjacent to excluded frames).
@@
 […]
-(a) surprised
-(b) disgusted
-(c) amused
-(d) scared
-(e) disappointed
-(f) satisfied
+(a) disappointed
+(b) satisfied
+(c) surprised
+(d) disgusted
+(e) amused
+(f) scared
 Figure 3: Reported emotions highly correlated with AU activation values.
-The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?”.
-Each subfigure shows results for a different emotion (indicated in the captions) and contains multiple pairs of boxplots representing the results for different AUs.
-The indices of the AUs are along the x-axis, while the y-axis represents the AU activation values in the reaction clip of each image.
-The parenthesized numbers in the legends give the number of responses of that type.
+The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?” Each subfigure shows the results for a different emotion (indicated in the subcaptions) and contains multiple pairs of boxplots representing the results for different AUs.
+The x-axis represents the indices of the AUs while the y-axis represents the AU activation values in the reaction clip of each image.
 In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion.
-In each boxplot, the bottom/top of the box represents the first/third quartile (25th/75th percentile) of the AU activation values, and the line in the middle of the box represents the median.
+In each boxplot, the bottom of the box represents the first quartile (25th percentile) of the AU activation values, the top of the box represents the third quartile (75th percentile), and the line in the middle of the box represents the median.
+The numbers of responses of each type for each subfigure are as follows.
+(a) Yes (592); No (1625).
+(b) Yes (724); No (1493).
+(c) Yes (352); No (1865).
+(d) Yes (124); No (2093).
+(e) Yes (368); No (1849).
+(f) Yes (194); No (2023).
 We started with analyzing statistical relationships between the AU activation values and answers in the annotation survey for each image. The primary results are as follows:
   • Overall Ratings.
-The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant positive correlation with the activation values of AU2 and AU12 and a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs.
-2(f), 2(a), 2(b) and 2(g) (Spearman correlation: $\rho\approx 0.09$ and $p<5\times 10^{-5}$ for AU2, $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9, $\rho\approx 0.08$ and $p<5\times 10^{-4}$ for AU12).
+The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant positive correlation with the activation values of AU2 and AU12 and a significant negative correlation with the activation values of AU4 and AU9 as shown in Figures 2(f), 2(a), 2(b) and 2(g) (Spearman correlation: $\rho\approx 0.09$ and $p<5\times 10^{-5}$ for AU2, $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9, $\rho\approx 0.08$ and $p<5\times 10^{-4}$ for AU12).
 […]
@@
 Each of the reported emotions was found to be significantly correlated with the activation values of multiple AUs.
+    • Disappointed.
+“Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Figure 3(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
+    • Satisfied.
+“Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Figure 3(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
     • Surprised.
-“Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig.
-3(a) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
+“Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Figure 3(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
     • Disgusted.
-“Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig.
-3(b) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
+“Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Figure 3(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
     • Amused.
-“Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig.
-3(c) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
+“Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Figure 3(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
     • Scared.
-“Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig.
-3(d) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
-    • Disappointed.
-“Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig.
-3(e) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
-    • Satisfied.
-“Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig.
-3(f) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
+“Scary” images are associated with higher activation values of AU4 and AU5 as shown in Figure 3(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
 Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$ (12 investigated AUs, 2 tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 emotions).
@@
 6.2.1 Pre-trained Scoring Models
-Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to estimate human preferences of text-to-image generation based on the input text prompts and the generated images.
+Researchers have used the CLIP score [60], Aesthetic score [65], BLIP score [43], ImageReward score [80], PickScore [35], and HPS v2 score [78] to estimate human preferences of text-to-image generation based on input text prompts and generated images.
 Among these baseline scoring models, ImageReward, PickScore, and HPS v2 were all trained specifically on datasets of human preferences of text-to-image generation and are supposed to outperform the other models [80, 35, 78].
@@
 To achieve this goal, we train an FAU-Net (Facial Action Units Neural Network) to do automatic scoring.
-In the FAU-Net, the estimated activation values of the 12 investigated AUs are used as the inputs; each activation value is first preprocessed with a linear transformation with trained parameters and sigmoid function ($\sigma(a\cdot\alpha+b)$ with the initializations of $a=1$ and $b=-1$ for encouraging transformations of positive correlation for better interpretability); then the preprocessed AU activation values are used as inputs to a hidden layer with 16 neurons which feeds into a single ouput neuron representing the AU evaluation score of the image; this output unit receives inputs from both the 16 neurons of the hidden layer and the 12 preprocessed AU activation values.
+In the FAU-Net, the estimated activation values of the 12 investigated AUs are used as the inputs; each activation value is first preprocessed with a linear transformation with trained parameters and sigmoid function ($\sigma(a\cdot\alpha+b)$ with the initializations of $a=1$ and $b=-1$ for encouraging transformations of positive correlation for better interpretability); then the preprocessed AU activation values are used as inputs to a hidden layer with 16 neurons which feeds into a single output neuron representing the AU evaluation score of the image; this output unit receives inputs from both the 16 neurons of the hidden layer and the 12 preprocessed AU activation values.
 The sigmoid function is used as the activation function for all neurons.
-The network is trained with a RankNet loss function using the reaction clips and ranking order given by the participant for image pairs generated from the same text prompts and an L1 regularization with a coefficient of $10^{-4}$.
-For model training, we use the Adam optimizer with an initial learning rate of $10^{-3}$ and a single batch for all data to train the neural network for 300 epochs using a single NVIDIA RTX A6000 48GB GDDR6 GPU for about 15 minutes.
+The network is trained with a RankNet loss function [5] using the reaction clips and ranking order given by the participant for image pairs generated from the same text prompts and an L1 regularization with a coefficient of $10^{-4}$.
+For model training, we use the Adam optimizer with an initial learning rate of $10^{-3}$ and a single batch for all data to train the neural network for 300 epochs using a single NVIDIA RTX A6000 48GB GDDR6 GPU for about 15 min.
 Our FAU-Net is trained with an LOPO procedure and outputs an FAU-Net valence score $s_{\mathrm{FAU}}$ reflecting the evaluation of the generated image derived from the facial expression reaction.
@@
 We separately investigated the effect of integrating the FAU-Net valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble.
-Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); the FAU-Net weights and the weightings $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, $w_{\mathrm{HPSv2}}$ computed in Eq.
-3 are not modified at this stage.
+Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); the FAU-Net weights and the weightings $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, $w_{\mathrm{HPSv2}}$ computed in Equation 3 are not modified at this stage.
 Table 2: Accuracy of image preference prediction for a total of 3734 image pairs.
@@
 Ensemble Baseline Score 65.80 68.64
+The results are shown in Table 2.
+Integration with FAU-Net valence score improves the performance of all four baseline models.
+Specifically, by further integrating all three pre-trained human preference scoring models together with the FAU-Net valence score, we achieve the highest accuracy of 68.64%, outperforming the ensemble of the three models by a margin of 2.84% ($p<0.01$ with a two-proportion z-test).
+6.2.4 Complementarity between FAU-Net and Pre-trained Scoring Models
+The complementarity between FAU-Net and pre-trained scoring models is demonstrated in Figure 4.
+The figure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (image pairs are selected based on the largest absolute difference scores for the two images given by the associated scoring model).
+We can clearly see that the three pre-trained scoring models are highly correlated showing similar performance curves (almost overlapping).
+In contrast, our FAU-Net valence score selects different image pairs, explaining why the FAU-Net valence scores combine synergistically with the other scores.
+To better understand the technical details, let’s first take Figure 4(d) as an example.
+In Figure 4(d), we first rank the image pairs in the dataset by the absolute value of the difference in FAU-Net scores for the two images, with the assumption that larger score difference implies higher confidence in predicting that the user prefers the image with higher score in the pair and thus higher accuracy.
+The x-axis of the figure represents “proportion selected”, which is the proportion of “most confident” image pairs—with left side meaning only the image pairs with highest confidence taken into account while right side meaning all image pairs taken into account.
+The red line in the figure represents the accuracies of FAU-Net for different proportions of “most confident” image pairs selected, and indeed as expected, the accuracy is clearly higher on the left side for the most confident image pairs, which confirms that our model is well calibrated.
+Now note that on the left side (low proportion of data selected) of the figure, the red line is obviously higher than the other lines, representing other pre-trained scoring models.
+This implies that for the image pairs with largest differences in their FAU-Net scores, the choice made by the FAU-Net is significantly more accurate than the choice made by the other models—in other words, for the image pairs our FAU-Net model is very good at, the other models are not good at them.
+At first glance, this might seem like a matter of course and not impressive, but this is actually not true for other models.
+Let’s then take Figure 4(a) as an example.
+As shown in its subcaption in Figure 4(a), the “most confident image pairs to compare” here are selected by choosing the image pairs with largest differences in their ImageReward scores.
+We can see that the ImageReward score (represented by the blue line) is well calibrated, but is also almost overlapping with the orange line and the green line—they are highly correlated.
+This implies that for the image pairs the ImageReward model is very good at, the PickScore and the HPS V2 models are also similarly very good at, so they are less complementary to each other.
+We can see similar results in Figures 4(c) and 4(b) as well.
+Overall in all of Figures 4(a), 4(c), 4(b) and 4(d), the curves for all three pre-trained scoring models are highly correlated while our FAU-Net score is not correlated with them.
+This demonstrates the uniqueness of our FAU-Net valence score, being very complementary to the other models.
+We can think about how it works as follows.
+For the image pairs with small differences in FAU-Net scores (for instance if a user does not show much facial expression difference), the FAU-Net scores will not affect the final score too much in Equation 5, so the other models dominate the decision.
+However, for the image pairs with large differences in FAU-Net scores, the FAU-Net scores dominate the final score in Equation 5, so our final choice decision is now more based on the FAU-Net model than the other models, and these are exactly the image pairs where our FAU-Net score has much better accuracy than the other models.
 (a) Selected by
@@
 (d) ${\color[rgb]{1,0,0}w_{h(4)-o}\approx-0.49}$
-(e) ${\color[rgb]{1,0,0}w_{h(5)-o}\approx-0.46}$
-(f) ${\color[rgb]{1,0,0}w_{h(6)-o}\approx-0.45}$
-(g) ${\color[rgb]{1,0,0}w_{h(7)-o}\approx-0.45}$
-(h) ${\color[rgb]{1,0,0}w_{h(8)-o}\approx-0.42}$
-(i) ${\color[rgb]{1,0,0}w_{h(9)-o}\approx-0.37}$
-(j) ${\color[rgb]{1,0,0}w_{h(10)-o}\approx-0.22}$
-(k) ${\color[rgb]{1,0,0}w_{h(11)-o}\approx-0.13}$
-(l) ${\color[rgb]{1,0,0}w_{h(12)-o}\approx-0.12}$
-(m) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
-(n) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
-(o) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
-(p) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
+(e) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
+(f) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
+(g) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
+(h) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
 Figure 5: Weights of hidden nodes.
-Each subfigure shows the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption illustrating the weight from the hidden node to the output node (red/green for negative/positive weights).
-The hidden nodes are ordered by the weights from them to the output node in this figure.
+Subfigures show the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption giving the weight from that hidden node to the output node (red/green for negative/positive weights).
+Four hidden nodes with lowest/highest negative/positive weights to the output node are shown here.
+Full version with all 16 hidden nodes is shown in the supplemental material.
 Figure 6: Weights of the output node. The weights from the input preprocessed activation values of the 12 AUs to the output node in the FAU-Net.
-The results are shown in Tab.
-2.
-Integration with FAU-Net valence score improves the performance of all four baseline models.
-Specifically, by further integrating all three pre-trained human preference scoring models together with the FAU-Net valence score, we achieve the highest accuracy of 68.64%, outperforming the ensemble of the three models by a margin of 2.84% ($p<0.01$ with a two-proportion z-test).
-6.2.4 Complementarity between FAU-Net and Pre-trained Scoring Models
-The complementarity between FAU-Net and pre-trained scoring models is demonstrated in Fig.
-4.
-The figure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (image pairs are selected based on the largest absolute difference scores for the two images given by the associated scoring model).
-We can clearly see that the three pre-trained scoring models are highly correlated showing similar performance curves (almost overlapping).
-In contrast, our FAU-Net valence score selects different image pairs, explaining why the FAU-Net valence scores combine synergistically with the other scores.
 6.2.5 FAU-Net Analysis
@@
 Generally, for hidden nodes positively affecting the output score, they receive positive weights from the activations of AU2, AU6, AU12, AU17, AU20, and AU26 and negative weights from the activations of AU4, AU5, AU9, AU17, and AU25, and for hidden nodes negatively affecting the output score, they receive weights of the opposite signs from the AUs.
-Similar trends are observed for the weights from the input preprocessed AU activation values and the output node as shown in Fig.
-6.
+Similar trends are observed for the weights directly from the input preprocessed AU activation values and the output node as shown in Figure 6.
 7 Discussion and Conclusion
@@
 We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
-Specifically, we developed an FAU-Net that receives the detected AU activations in the facial expression reaction as inputs and outputs an FAU-Net valence score as the estimation of the user preference.
+Specifically, we developed an FAU-Net that receives the detected AU activations in the facial expression reaction as inputs and outputs an FAU-Net valence score as an estimation of user preference.
 Integrating the FAU-Net valence score with pre-trained scoring models improves their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
-Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al.
-[79] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.47% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4.
+Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: [80] used ImageReward (trained from manually labeled human preferences), with only an accuracy of 65.14% on their own dataset and 60.47% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4.
 (see Tab.
@@
 6 of [80]).
-We also demonstrated the complementarity between our FAU-Net and the pre-trained scoring models, explaining why integrating them effectively improves the annotation accuracy.
-As they are respectively better at annotating different sets of image pairs, it is also potentially feasible to use them to fine-tune each other on data where one model gives larger absolute difference of scores for the two images.
+We also demonstrated the complementarity between FAU-Net and the pre-trained scoring models, explaining why integrating them effectively improves the annotation accuracy.
+Admittedly, our study has some limitations: (1) users may be reluctant to turn on their cameras in real life; (2) the participants of our study are explicitly aware that their reaction videos are being recorded, which might affect the spontaneousness of their facial expressions [87, 55].
 In conclusion, we have demonstrated the feasibility of automatically scoring user preferences for image generation from facial expression reaction.
-The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [57], image inpainting [22], and super-resolution [74].
+The application of this method may also be potentially generalizable to other image generation tasks, such as image-to-image translation [58], image inpainting [22], and super-resolution [75].
 8 Acknowledgments
-We thank Xiaojing Xu and Yuan Tang for helpful prior work and Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions.
-We are grateful for support from NSF IIS 1817226 and seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion as well as hardware funding from NVIDIA, Adobe, and Sony.
+We thank Xiaojing Xu and Yuan Tang for helpful prior work, Gary Cottrell, Vijay Veerabadran, and Miguel Monares for helpful discussions, and reviewers for helpful comments.
+We are grateful for support from NSF IIS 1817226 and CRCNS 2208362, seed funding from UC San Diego Social Sciences and the Sanford Institute for Empathy and Compassion, and gift support from an anonymous donor, as well as hardware funding from NVIDIA, Adobe, and Sony.
 Supplementary Material
 A Supplementary Visuals
-To better illustrate the idea of our paper, we present Fig.
-S1 showing an example of a strong AU4 activation in response to a low-quality image generation and Fig.
-S2 an example of a strong AU12 activation in response to a high-quality image generation.
+To better illustrate the idea of our paper, we present Figure S1 showing an example of a strong AU4 activation in response to a low-quality image generation, and Figure S2 an example of a strong AU12 activation in response to a high-quality image generation.
 Figure S1: An example of AU4 activation in response to a low-quality image generation.
@@
 B Data Collection
-This section provides more details of the collection of the FERGI dataset not provided in Sec.
-3.
-Figs.
-S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
+Figure S3: Procedure of one session in data collection.
+A copy of Figure 1 with additional references to the screenshots of each stage.
+Firstly, the participant creates an input text prompt.
+Secondly, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately; at the same time, 5 images are generated from the text prompt using Stable Diffusion v1.4.
+Then, for each image, the participant goes through a 5-second baseline phase, a 5-second image presentation phase, and an image annotation phase with no time restriction.
+(During the baseline phase, the participant’s baseline facial expression is recorded, and during the image presentation phase, the participant’s facial expression reaction to the generated image is recorded.) Finally, after all 5 images have been presented and annotated, the participant ranks the 5 generated images from best to worst.
+This section provides more details of the collection of the FERGI dataset not provided in Section 3.
+Figure S3 displays a copy of Figure 1, showing the procedure of one session in data collection with additional references to the screenshots of each stage.
+Figures S4, S5, S6, S7, S8, S9 and S10 display screenshots of an example of one session in data collection.
 B.1 Participants
-39 participants [^1] The user IDs are up to 040 as shown in Tab.
-S1 because the same participant used both 027 and 028 due to some technical issues.
+39 participants [^1] The user IDs are up to 040 as shown in Table S1 because the same participant used both 027 and 028 due to some technical issues.
 Though, no data was collected under the user ID 027.
@@
 Most participants completed the same number of structured input sessions and free-form input sessions.
-The number of sessions completed by each participant are shown in Tab.
-S1.
+The number of sessions completed by each participant are shown in Table S1.
 By default, each participant is expected to complete 9 structured input sessions and 9 free-form input sessions (for the first 13 out of the 33 participants chronologically) or to complete 10 structured input sessions and 10 free-form input sessions (for the last 20 out of the 33 participants chronologically).
@@
 internet connection), and various technical issues also occur occasionally, so the total number and categories of completed sessions also vary.
+Table S1: Information of the FERGI dataset. Numbers of structured input (SI) sessions, numbers of free-form input (FFI) sessions, and numbers of valid images for each participant are shown.
 User ID Number of SI Sessions Number of FFI Sessions Number of Valid Images
@@
 Total 288 288 2827
-Table S1: Information of the FERGI dataset. Numbers of structured input (SI) sessions, numbers of free-form input (FFI) sessions, and numbers of valid images for each participant are shown.
 B.2.1 Structured Input
@@
 Structured input contains 5 sections:
-  • Animate Objects (Fig.
-3(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc.
+  • Animate Objects (Figure 4(a)): Any object that can do things, including humans, animals, robots, fantasy creatures, etc.
 Participants are required to enter the name and quantity of the object, with activity and a list of characteristics as optional inputs.
-  • Inanimate Objects (Fig.
-3(b)): Any object that cannot do things, including plants, vehicles, furniture, etc.
+  • Inanimate Objects (Figure 4(b)): Any object that cannot do things, including plants, vehicles, furniture, etc.
 Participants are required to enter the name and quantity of the object, with a list of characteristics as optional inputs.
-  • Interactional Relations (Fig.
-3(c)): Relationship between an animated object and any other object (something an animate object does to another object).
+  • Interactional Relations (Figure 4(c)): Relationship between an animated object and any other object (something an animate object does to another object).
 An animate object can only be in a relation if its “activity” field is empty.
 All objects can only be involved in at most one relation (either interactional or positional).
-  • Positional Relations (Fig.
-3(d)): Relationship between any two objects.
+  • Positional Relations (Figure 4(d)): Relationship between any two objects.
 An animate object can only be in a relation if its “activity” field is empty.
 All objects can only be involved in at most one relation (either interactional or positional).
-  • Other Inputs (Fig. S4): The location of image as background or general environment, the style of the image, and a list of keywords to append at the end of the prompt.
+  • Other Inputs (Figure S5): The location of image as background or general environment, the style of the image, and a list of keywords to append at the end of the prompt.
 After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model).
 Participants can freely adjust the generated prompt as the final prompt.
-Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs.
-S3 and S4.
+Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figures S4 and S5.
 B.2.2 Free-Form Input
@@
 The participants are provided with an example prompt and a text box to enter the entire prompt as the final prompt.
-A screenshot of an example of the prompt creation page for a free-form input session is shown in Fig.
-S5.
+A screenshot of an example of the prompt creation page for a free-form input session is shown in Figure S6.
 B.3 Image Presentation
-Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
+Screenshots of an example of the pages for image presentation and its baseline are shown in Figures S7 and S8.
 B.3.1 Image Annotation
@@
   • Participants can optionally select from a list of emotions to report their feelings when seeing the image.
-A screenshot of an example of the image annotation page (including the details of the questions) is shown in Fig. S8.
+A screenshot of an example of the image annotation page (including the details of the questions) is shown in Figure S9.
 B.3.2 Image Ranking
@@
 All 5 images are displayed side by side, with the left side labeled “Best” and right side labeled “Worst”, and the participants provide ranking by dragging the images to reorder them.
-A screenshot of an example of the image ranking page is shown in Fig.
-S9.
+A screenshot of an example of the image ranking page is shown in Figure S10.
 (a) Animate objects
@@
 C AU Model Training
-This section provides more details of the training of the AU model not provided in Sec. 4.
+This section provides more details of the training of the AU model not provided in Section 4.
 C.1 Datasets
@@
 Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5.
-See Fig.
-S10 for a visual reference guide for the analyzed AUs.
+See Figure S11 for a visual reference guide for the analyzed AUs.
 Figure S11: A visual reference guide for the analyzed AUs extracted from [57].
@@
 $CE(y,\hat{y})=-[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})].$ (10)
-The weights for the MSE loss and those for the cross entropy loss are both inverse-frequency weighted (for the MSE loss these are based on the merged groups of $\{0,1\}$ and $\{2,3,4,5\}$).
+The weights for the MSE loss and those for the cross entropy loss are both inverse-frequency weighted and normalized within each AU for addressing class imbalance in the datasets (substantially higher number of occurrences for low AU intensities).
+However, since the number of occurrences of high intensities are too few for most AUs (resulting in too high weights for the MSE loss if used directly), we “bin” the intensities into 2 groups, and each group shares the same weight.
+Specifically, for the MSE loss, we apply one weight for occurrences of intensities of 0 and 1 and another weight for occurrences of intensities of 2, 3, 4, and 5, and these weights are computed based on the total number of occurrences within each intensity group.
 Specifically, the weights for the MSE loss are defined as
@@
 D Data Filtering
-This section provides the details of frame exclusion not provided in Sec. 5.1.
+This section provides the details of frame exclusion not provided in Section 5.1.
 We aim to exclude frames with occlusions and frames with off-angle pitch or yaw. To achieve the goal, we follow the following three exclusion rules:
@@
     where $d_{\textrm{nostrils-eyes}}$ represents the vertical distance between the center of the nostrils and the center of the eyes while $d_{\textrm{eyes}}$ represents the horizontal distance between the two eyes, both of which are computed based on the facial landmarks detected with [8].
+E Experiments
+E.1 FAU-Net Analysis
+A full version of Figure 5 showing the weights of the hidden nodes in the FAU-Net is shown as Figure S12.
+(a) ${\color[rgb]{1,0,0}w_{h(1)-o}\approx-0.58}$
+(b) ${\color[rgb]{1,0,0}w_{h(2)-o}\approx-0.53}$
+(c) ${\color[rgb]{1,0,0}w_{h(3)-o}\approx-0.50}$
+(d) ${\color[rgb]{1,0,0}w_{h(4)-o}\approx-0.49}$
+(e) ${\color[rgb]{1,0,0}w_{h(5)-o}\approx-0.46}$
+(f) ${\color[rgb]{1,0,0}w_{h(6)-o}\approx-0.45}$
+(g) ${\color[rgb]{1,0,0}w_{h(7)-o}\approx-0.45}$
+(h) ${\color[rgb]{1,0,0}w_{h(8)-o}\approx-0.42}$
+(i) ${\color[rgb]{1,0,0}w_{h(9)-o}\approx-0.37}$
+(j) ${\color[rgb]{1,0,0}w_{h(10)-o}\approx-0.22}$
+(k) ${\color[rgb]{1,0,0}w_{h(11)-o}\approx-0.13}$
+(l) ${\color[rgb]{1,0,0}w_{h(12)-o}\approx-0.12}$
+(m) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
+(n) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
+(o) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
+(p) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
+Figure S12: Weights of hidden nodes.
+Each subfigure shows the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption illustrating the weight from the hidden node to the output node (red/green for negative/positive weights).
+The hidden nodes are ordered by the weights from them to the output node in this figure.